### v3

# Jingjing Sun, Pengyi Shi, Jim Dai

[…]
$0$
[…]
Jingjing Sun, Pengyi Shi, Jim Dai
[…]
Inpatient Overflow Management with Proximal Policy Optimization
[…]
Inpatient Overflow Management
with Proximal Policy Optimization
[…]
Problem Definition: Managing inpatient flow in large hospital systems is challenging due to the complexity of assigning randomly arriving patients – either waiting for primary units or being overflowed to alternative units. Current practices rely on ad-hoc rules, while prior analytical approaches struggle with the intractably large state and action spaces inherent in patient-unit matching. A scalable decision-support framework is needed to optimize overflow management while accounting for time-periodic fluctuations in patient flow. Methodology/Results: We develop a scalable decision-making framework using Proximal Policy Optimization (PPO) to optimize overflow decisions in a time-periodic, long-run average cost setting. To address the combinatorial complexity, we introduce atomic actions, which decompose multi-patient routing into sequential, tractable assignments. We further enhance computational efficiency through a partially-shared policy network designed to balance parameter sharing with time-specific policy adaptations, and a queueing-informed value function approximation to improve policy evaluation. Our method significantly reduces the need for extensive simulation data, a common limitation in reinforcement learning applications. Case studies on hospital systems with up to twenty patient classes and twenty wards demonstrate that our approach matches or outperforms existing benchmarks, including approximate dynamic programming, which is computationally infeasible beyond five wards. Managerial Implications: Our framework offers a scalable, efficient, and adaptive solution for managing patient flow in complex hospital systems. More broadly, our results highlight that domain-aware adaptation – leveraging queueing structures and operational insights – is more critical to improving algorithm performance than fine-tuning neural network parameters when applying general-purpose algorithms to specific domain applications.
[…]
1 Introduction
Managing inpatient flow is a significant challenge in hospital operations, especially when patient demand fluctuates unpredictably. Hospital capacity often cannot be scaled up quickly enough when demand increases, resulting in intermittent bed shortages and overcrowding. Overcrowded inpatient units can lead to a range of problems, including denied admissions, emergency department (ED) boarding, ambulance diversions, surgery cancellations, and adverse patient outcomes caused by long wait times (McKenna et al. 2019, Jones et al. 2022). The issue of intermittent bed shortages is worsened by the common practice of grouping specific clinical services within designated inpatient units, each with a fixed number of beds (Song et al. 2020). When the designated inpatient units of certain clinical services reach capacity, hospitals often have to find alternative ways for their incoming patients, which frequently involves assigning them to non-dedicated “overflow” units.
This overflow practice, also known as off-service placement (Song et al. 2020) or patient outlying (Sprivulis et al. 2006), helps temporarily alleviate congestion in hospitals by assigning patients to non-primary units when their designated wards are full. This practice is commonly used worldwide, including in the United States (Zhang et al. 2020, Lim et al. 2024), England (Stylianou et al. 2017), France (Stowell et al. 2013), Spain (Alameda and Suárez 2009), Australia (Sprivulis et al. 2006), and Singapore (Shi et al. 2016), with 10% to 20% of patients placed in overflow units according to these studies. However, this practice introduces additional complexity into patient placement and may compromise the quality and continuity of care (Stowell et al. 2013, Stylianou et al. 2017, Song et al. 2020). Moreover, excessive reliance on overflow can reduce overall system efficiency by increasing coordination efforts and potentially triggering a cycle of subsequent overflow events, further exacerbating congestion (Dong et al. 2019).
Effectively managing overflow requires nuanced decision-making to balance the benefits of reducing inpatient overcrowding with the costs of unnecessary or excessive overflow assignments. It is crucial to account for the long-term impact, given that the typical patient length-of-stay is 3 to 4 days. This means that an overflow decision not only affects the immediate system state but can also influence system congestion over several days. This complexity is further compounded by time-varying dynamics, such as predictable hourly patterns of patient arrivals and discharges. For example, should patients wait a few more hours if more discharges are expected soon? Overflow assignments in practice, though widely used, often rely on anecdotal experiences or ad-hoc rules, which are suboptimal and lack the rigorous analytical support needed for effective decision-making. Dai and Shi (2019) is one of the earliest works that formulate the overflow decision problem as a periodic, long-run average-cost Markov Decision Process (MDP), aiding the decisions on whether to overflow a waiting patient and, if so, to which ward. Several other recent works have also adopted analytical methods to aid overflow decision making (Schmidt et al. 2013, Zhang et al. 2020, Bertsimas and Pauphilet 2024, Dean 2024).
These analytical approaches, however, share a common challenge: the intractably large state and action spaces inherent in the “matching problems.” In the context of patient overflow, this arises from multiple patients (customers) from different specialties waiting for admission, while a combinatorial number of bed assignment (server) options – including the choice to continue waiting – are available. Similar challenges are observed in other domains like ride-hailing, inventory management, and volunteer assignment, where the combinatorial nature of potential actions greatly complicates decision-making. Dai and Shi (2019) tackle the large state space using value function approximation but encounter difficulties with the large action space, which limited their method’s scalability – their case study only dealt with systems involving up to five specialties and wards. Other approaches, such as Linear Programming or fluid-based optimization, manage large action spaces more efficiently but often oversimplify future stochastic variations (Bertsimas and Pauphilet 2024); see a comprehensive literature review in Section 2. These limitations highlight the broader difficulty in developing practical and scalable decision-support algorithms for complex customer-server matching problems, not just to healthcare but a wide range of different domains.
1.1 Paper Overview and Contribution
This paper addresses the existing gap by developing an algorithmic framework based on the Proximal Policy Optimization (PPO) method to minimize the long-run average costs associated with overflow. Our key contributions are summarized as follows.
  1. 1.
    Scalable Algorithim: We develop the first known scalable algorithm for overflow decision-making in a time-periodic, long-run average setting. Our algorithm successfully handles systems with up to 20 medical specialties and 20 hospital units, greatly exceeding the scale of prior methods. Beyond healthcare, our framework is applicable to a broad class of matching problems in operations management with similar queueing structures and contributes to efficiently managing the combinatorial complexity in customer-server matching across domains.
    The success of this framework hinges on two key elements: (a) leveraging the “atomizable” action space and (b) incorporating domain knowledge of queueing systems – particularly in designing value function approximation and policy representation. We elaborate on them next.
  2. 2.
    Action Atomization: We propose the atomic action in Section 4.1, which decomposes class-level decisions into sequential individual-level assignments. This transformation fundamentally simplifies the original combinatorial action space and allows for efficient policy optimization via gradient-based methods. Despite its seemingly myopic nature, we demonstrate in our case study that this approach achieves performance comparable to brute-force optimization over the entire action space (Dai and Shi 2019), i.e., this approach works. Moreover, it retains flexibility for aggregate planning and incorporating individual preferences for practical implementation.
  3. 3.
    Dealing with Time-periodicity: To account for real-world daily and weekly periodic patterns in patient arrivals and discharges, we integrate time-periodicity into our PPO framework by leveraging the queueing system structure. First, we establish a theoretical guarantee for policy improvement of PPO under the periodic setting (Section 4.2). Second, we design a partially-shared policy network that balances parameter sharing and temporal specificity (Section 5.1). This design captures structural similarities while being adaptive to time variations. It outperforms two conventional alternatives: (i) a fully-shared network that requires more complex policy networks to model time dependencies, and (ii) a fully-separate network that demands significantly more training data. Lastly, we propose a new value function approximation derived from the Poisson equation and tailored to the queueing dynamics under the atomic actions – it decomposes the system-wide value function into pool-wise approximations to improve computational efficiency and stability (Section 5.2). Case study results in Section 6 confirm that our framework, which integrates all these design elements, enhances sample efficiency while maintaining superior performance over other benchmarks.
2 Literature review
2.1 Bed Partition and Patient Routing
Adding new beds to alleviate inpatient congestion is often not feasible due to constraints related to space, finances, and regulations. Therefore, we focus on literature that explores strategies for utilizing existing bed resources more efficiently. Best et al. (2015) examine bed management at a strategic level, where the decision involves partitioning inpatient beds into separate wings dedicated to different specialties (overflow is not allowed). As a follow-up, Izady and Mohamed (2021) consider a similar setting but introduce the option of “pooled wings,” which can accommodate patients from multiple specialties when dedicated wings reach capacity. Optimization techniques are primarily employed in this type of strategic planning to maximize inpatient bed utilization.
Song et al. (2020) highlight that wing designs can exacerbate inpatient crowding, particularly when combined with unpredictable patient volumes and limited bed capacity. This often results in frequent off-service (overflow) placements, especially in large teaching or public hospitals that handle high volumes of unpredictable emergency department demand (Manning and Islam 2023). However, it is crucial to manage these overflow assignments carefully to minimize adverse patient outcomes and ensure quality care (Lim et al. 2024). Regarding patient assignment under fixed bed partitioning, the most relevant study is Dai and Shi (2019), which serves as the foundation for this paper. They formulate the problem as an MDP and use approximate dynamic programming (ADP) to address the large state space. Their approach focuses on a deterministic policy space, which can handle systems with up to five server-pools but does not address the large action space. Chen et al. (2021) develop a primal-dual method and use a similar MDP for inpatient assignments as one of their case studies. Their method tackles the large state space by decomposing the problem into multiple subproblems with an inverted-V structure (i.e., single-class multi-pool). However, the action space in each subproblem remains combinatorial. In contrast, the PPO algorithm we propose in this paper addresses both the large state and action spaces and can effectively solve problems in large systems (e.g., 20 server pools). Dong et al. (2025) use ADP to optimize scheduling policies in multiclass, multiserver queues with wait-dependent service slowdowns, focusing on leveraging the problem structure to estimate value function differences directly rather than addressing action space complexity. Other recent studies applying ADP methods to patient-provider matching in healthcare operations include Cire and Diamant (2022), Gao et al. (2024), Khorasanian et al. (2024), among others.
In contrast to studying the routing problem from a sequential (dynamic) decision-making perspective, Thomas et al. (2013) formulate the patient assignment problem as a mixed-integer programming (MIP) model. Their objective is to minimize immediate costs without considering the long-term impact of current actions, which results in a myopic policy. Similarly, Bertsimas and Pauphilet (2024) propose a robust multistage optimization framework that minimizes the current cost along with an approximation of future costs based on a deterministic model. While their approach provides flexibility for individualized policies, it follows a different methodological framework. In particular, future costs are handled differently – they are approximated as a weighted sum over the remainder of the day and the following week, based on the worst-case scenarios within a predefined uncertainty set. This can lead to more conservative, risk-averse policies. In contrast, our approach directly optimizes policy decisions under uncertainty via PPO, focusing on the stochastic nature of the problem. Additionally, this paper emphasizes the importance of tailoring a general-purpose PPO algorithm to the structure of the inpatient overflow problem (e.g., specialized policy NN designs and basis functions) and demonstrates that these modifications can have a greater impact on performance than hyperparameter tuning.
2.2 PPO Applications
The PPO algorithm was first introduced in the seminal work by Schulman et al. (2017), which provided theoretical justification for using PPO to solve MDP problems with bounded cost functions and discounted cost objectives. Dai and Gluzman (2022) extend these theoretical guarantees to MDP problems with unbounded cost functions and long-run average cost objectives. Since then, numerous studies have successfully applied PPO to domains such as inventory management (Perez et al. 2021) and transshipment (Vanvuchelen et al. 2024). However, our inpatient overflow MDP problem presents unique challenges that make the direct application of PPO algorithms difficult, specifically due to the large action space and the time-varying, periodic long-run average setting.
In conventional PPO algorithms, the policy space is directly parameterized by an NN, with the output size corresponding to the number of possible actions – typically small in other applications (e.g., Schulman et al. (2017) and Dai and Gluzman (2022) focus on smaller action spaces). However, in our problem, the action space is significantly larger, making it computationally prohibitive to train the policy network using traditional parameterization methods. To our knowledge, there are no established examples in the literature where PPO has been successfully applied to such large action spaces, except for Feng et al. (2021), which showcase an effective use of action decomposition in ride-hailing applications. However, their setting is finite-horizon, whereas the framework we develop in this paper tackles the more complex infinite-horizon problem. Additionally, to account for the periodic nature of our infinite-horizon MDP, we introduce a novel partially-shared network structure that outputs actions at different epochs throughout the day. This design is conceptually related to specialized architectures in federated learning (Arivazhagan et al. 2019, Kulkarni et al. 2020), which address statistical heterogeneity, a different problem from ours.
3 Model Formulation
In this section, we present the queueing model formulation abstracted from the overflow problem. While we describe a general queueing model, we ensure it retains the key features relevant to our application. Specifically, the model is based on a parallel server system with multiple customer classes and server pools. Each customer class represents a medical specialty (or a cluster of related specialties), and each server pool corresponds to an inpatient unit (ward), with servers representing individual inpatient beds. We detail the queueing discipline along with the arrival and departure processes in Section 3.1. We formulate the overflow decision problem as an MDP in Section 3.2. Table 5 in Appendix 8 summarizes all the notations introduced in the model formulations and those used in later sections.
3.1 Parallel-server System
Consider a queueing system with $J$ customer classes and $J$ parallel-server pools. Each pool has $N_{j}$ identical servers and is dedicated to serving one of the $J$ customer classes – the primary pool for that class. Without loss of generality, we assume that server-pool $j$ is the primary pool for class $j$ customers, $j\in\mathcal{J}=\{1,2,\dots,J\}$. Upon the arrival of a class $j$ customer, if there are idle servers in primary pool $j$, the customer is admitted immediately, known as a primary assignment. Otherwise, the customer may wait in buffer $j$ (which has no capacity limit) for service in the primary pool or, at a decision epoch, be assigned to a non-primary pool according to the overflow policy specified in Section 3.2. This latter scenario is called an overflow assignment, and the non-primary pool that can serve class $j$ customers is referred to as an overflow pool for class $j$.
[…]
Departure process. We adopt a two-time-scale departure scheme as in Dai and Shi (2019). For customers admitted to pool $j$, there is a probability $\mu_{j}$ that they will depart from the system on the current day, and a probability $1-\mu_{j}$ that they will stay (memoryless). If a customer departs, the time of departure within the current day is modeled by the random variable $h_{\textrm{dis}}$, which follows a general distribution with support on $[0,1)$ and cumulative distribution function (CDF) $F_{j}$. This scheme was first introduced in Shi et al. (2016) and has since been widely used in modeling inpatient flow, e.g., Dai and Shi (2019), Bertsimas and Pauphilet (2024), Kim et al. (2015).
3.2 MDP Model
We use an MDP framework to decide on whether to assign a waiting customer to one of the overflow pools. The objective is to find the optimal overflow policy that minimizes the infinite-horizon, long-run average cost. We consider a discrete-time setup, where the decision maker (e.g., hospital manager) makes overflow decisions at $m$ predefined time epochs each day (e.g., hourly), denoted as $t_{0},t_{1},\dots,t_{m-1}$ for day $t$. In the rest of this section, we define the state, action, cost, policy, and objective of the MDP.
State.
We follow Dai and Shi (2019) to define the state space. Denote the system state observed at decision epoch $t_{k}$, before taking any action, as $S(t_{k})$, where
$S(t_{k})=\big{(}X_{1}(t_{k}),\dots,X_{J}(t_{k}),Y_{1}(t_{k}),\dots,Y_{J}(t_{k}),h(t_{k})\big{)}.$
The state variable is a $(2J+1)$-dimensional vector, with each component as follows:
  • $X_{j}(t_{k})$, $j=1,\dots,J$, denotes the customer count for pool $j$. This customer count includes both the number of customers waiting in buffer $j$ and those in service in pool $j$; we show below on how to recover each part. The former part (waiting count) only includes class $j$ customers, while the latter (in-service count) may include customers from other classes due to overflow assignments.
  • $Y_{j}(t_{k})$, $j=1,\dots,J$, denotes the to-depart count in pool $j$. This quantity represents the number of customers expected to depart from pool $j$ between the present time and the end of the current day. Tracking this quantity is essential for the two-time-scale inpatient departure scheme, where the “planned discharge” information is used to inform overflow decisions.
  • $h(t_{k})=k\in\{0,1,2,\dots,m-1\}$ denotes the time-of-day indicator.
We denote the state space as $\mathcal{S}$, and the sub-space that contains all states with time-of-day being $h$ as $\mathcal{S}^{h}$. Therefore, $(\mathcal{S}^{0},\dots,\mathcal{S}^{m-1})$ is a partition for $\mathcal{S}$ without overlap between sub-spaces.
Action.
We assume that all primary assignments are made immediately when customers arrive or when primary servers become available. Thus, at each epoch $t_{k}$, the decision maker only needs to decide whether to assign a waiting customer to an overflow pool or let them continue waiting. We use $f(t_{k})=\{f_{i,j}(t_{k}),~{}i,j=1,\dots,J\}$ to represent the overflow/wait decision at time $t_{k}$. Each $f_{i,j}$ corresponds to the number of assignments from class $i$ to pool $j$. If $i\neq j$, it denotes the number of overflow assignments; if $i=j$, it denotes the number of customers who remain waiting.
For a given state $S(t_{k})=s$, action $f$ is called feasible if it satisfies the following conditions:
$\sum_{\ell=1,\ell\neq j}^{J}f_{\ell,j}\leq(N_{j}-x_{j})\vee 0,\quad\sum_{\ell=1}^{J}f_{i,\ell}=(x_{i}-N_{i})\vee 0,\quad i,j=1,\dots,J.$ (1)
To explain, the total number of overflow assignments from all non-primary classes to pool $j$ must not exceed the number of idle servers in pool $j$, which can be computed as $(N_{j}-x_{j})\vee 0$, with $x\vee 0=\max(x,0)$ for any real number $x$. Similarly, the total number of assignments for class $i$, including overflow assignments to a pool $\ell\neq i$ and those kept waiting, must equal the number of waiting customers. This can be calculated as $(x_{i}-N_{i})\vee 0$, which we also refer to as the queue length. We use $q_{i}=(x_{i}-N_{i})\vee 0$ to denote this queue length for class $i$ (buffer $i$).
State transition.
We denote the state at epoch $t_{k}$ before taking an action as $s$, i.e., the pre-action state. After an action is taken, the resulting state is denoted as $s^{+}$, i.e., the post-action state. Given an action $f={f_{i,j}}$, the transition from $s$ to $s^{+}$ is as follows:
$x^{+}_{j}$ $=x_{j}+\sum_{i=1,i\neq j}^{J}f_{i,j}-\sum_{\ell=1,\ell\neq j}^{J}f_{j,\ell},\quad j=1,\dots,J;$ (2)
$y_{j}^{+}$ $=y_{j},\quad j=1,\dots,J;$
$h^{+}$ $=h.$
We denote the state, after the exogenous events (random arrivals and departures) occurring between the current epoch $t_{k}$ and the next epoch $t_{k+1}$, as $s^{\prime}$. The corresponding transition probability from state $s$ to state $s^{\prime}$, through $s^{+}$ under action $f$, is denoted as $p(s^{\prime}|s,f)$. The detailed calculation of these transition probabilities is provided in Appendix 9.
Cost.
As discussed, there are negative consequences associated with both excessive waiting (system congestion) and overflow assignments. To capture this trade-off, we consider two types of costs: (i) holding cost: each waiting customer of class $j$ incurs a cost $C_{j}$ at each epoch; (ii) overflow cost: each overflow assignment from class $i$ to pool $j$ incurs a one-time cost $B_{i,j}$. Given pre-action state $s$ and action $f$, the one-epoch cost is the sum of these two components, expressed as:
$g(s,f)=\sum_{j=1}^{J}C_{j}q_{j}^{+}+\sum_{i=1}^{J}\sum_{j=1,j\neq i}^{J}B_{i,j}f_{i,j},$
where $q_{j}^{+}=(x_{j}^{+}-N_{j})\vee 0$ denotes the post-action queue length for class $j$, and the transition to the post-action state follows (2). This cost setup is designed to identify a policy that strikes a balance between the negative effects of waiting and the undesirability of overflow assignments.
Policy and objective.
We consider the long-run average cost objective since overflow decisions have long-term impacts on the system. We focus on all feasible stationary Markovian policies, and we denote the policy space as $\Pi$. For a policy $\pi\in\Pi$, we use $\pi(f|s)$ to denote the probability of selecting a feasible action $f$ in a given state $s$ (where deterministic policies are included via using the Dirac measure). The resulting stochastic process $\{S(t_{k})\}$ forms a discrete-time Markov Chain (DTMC) with a one-epoch transition probability from state $s$ to $s^{\prime}$ as $p_{\pi}(s^{\prime}|s)=\mathbb{E}_{f\sim\pi(\cdot|s)}p(s^{\prime}|s,f)$. The expected one-epoch cost is $g_{\pi}(s)=\mathbb{E}_{f\sim\pi(\cdot|s)}g(s,f)$. Correspondingly, the objective is to identify the optimal policy
$\pi^{*}=\arg\min_{\pi\in\Pi}\lim_{T\rightarrow\infty}\frac{1}{T}\mathbb{E}_{\pi}\left[\sum_{t=1}^{T}\sum_{k=0}^{m-1}g_{\pi}\big{(}S(t_{k})\big{)}\right].$ (3)
Under the periodic arrival assumptions, the system reaches a periodic steady state (Liu and Whitt 2011a, b). Without loss of generality, we assume that there exists a stationary Markovian policy that is optimal; see Puterman (2014) for conditions under which this holds. We conclude this section with the following remarks on possible extensions.
Extensions. We can use the concise state representation with customer count $X_{j}$ because we assume (i) primary assignments are handled automatically on-the-fly, allowing us to calculate both the waiting count and the in-service count from $X_{j}$, and (ii) the departures depend only on the server pool, so we can disregard customer class information once a customer is admitted into service. For (i), our analytical framework can easily be extended to a bed assignment problem that incorporates primary assignments (not assumed to be handled automatically). This would require augmenting the state space by separating $X_{j}$ into the waiting count $Q_{j}$ and the in-service count $Z_{j}$, where $Q_{j}$ tracks all customers who arrived between decision epochs, including those who could have been assigned to primary pools upon arrival but are now waiting. Assumption (ii) is made due to the relatively consistent average length of stay across medical specialties (approximately 4 days), while it could be relaxed in a similar way by augmenting the state space.
While our baseline setting does not consider transferring overflow customers back to their primary pool, this can be accommodated by adding actions that represent the number of in-service transfers from each pool $j$ back to their primary pools. In addition, although we assume a linear cost function in this paper, our framework can handle more complex cost structures, such as polynomial costs or costs related to service levels (e.g., the proportion of patients waiting more than 6 hours). See the appendix of Dai and Shi (2019) for an example of modifying the cost to account for service-level metrics. We note that these adjustments do not alter the core elements of our problem, and our algorithmic framework is both flexible and general enough to accommodate them. For ease of exposition, throughout the remainder of the paper, we will focus on the current baseline setting, where primary assignments are handled automatically, departures are pool-dependent, no transfer back of overflow customers, and the cost is linear.
4 Atomic Actions and Proximal Policy Optimization
One of the core challenges in solving the MDP model developed in Section 3.2 is the combinatorial explosion of the action space. When assigning servers for multiple customers simultaneously, the number of possible combinations grows exponentially with the number of customer classes and server pools. Specifically, when there are enough idle servers in the system, the action space has a size of $\prod_{j=1}^{J}\left(\begin{array}[]{c}q_{j}+w_{j}\\ w_{j}\end{array}\right)$, where $q_{j}$ is the number of waiting customers in class $j$ and $w_{j}$ is the number of overflow pools for class $j$ that have idle servers. To illustrate, in a twenty-pool system with each server-pool having a capacity of 60 and a state $(x_{1},x_{2},x_{3},\dots,x_{20})=(65,63,62,50,50,50,\dots,50)$, the number of potential ways to assign 10 waiting customers can exceed $10^{9}$. In other words, even if the value function is computed (either exactly or via approximation), brute-force searching through all possible actions to find the one with the lowest cost-to-go is infeasible. To overcome this curse of dimensionality, we use atomic actions in combination with PPO to directly optimize the policy, which forms the key foundation for our algorithm’s scalability. In Section 4.1, we introduce the atomic action and policy setup. In Section 4.2, we present the framework of the PPO algorithm.
4.1 Action Space Reduction: Atomic Actions
Atomic actions break the original action into a sequence of individual ones. That is, instead of solving the assignments for all waiting customers at once, the atomic action approach decomposes the problem into smaller, individual actions, sequentially making overflow/wait decisions for each customer. This sequential decision-making process is referred to as the atomic decision process.
To elaborate, for a given state $s=(x_{1},\dots,x_{J},y_{1},\dots,y_{J},h)$ observed at epoch $t_{k}$, the total number of waiting customers across all classes is $q=\sum_{j=1}^{J}q_{j}$, where $q_{j}=(x_{j}-N_{j})\vee 0$. Following a pre-determined rule (e.g., first-come-first-serve), we determine an order for these $q$ waiting customers and conduct atomic actions $\mathbf{a}=\{a^{1},a^{2},\dots,a^{q}\}$ for each of them sequentially. Here, $a^{n}$ represents the assignment for the $n$th customer in the chosen order. The action space for each $a^{n}$ is simply $\mathcal{J}=\{1,\dots,J\}$. If the class of the $n$th customer is $j$, then $a^{n}=j$ indicates the customer will continue waiting until the next decision epoch; otherwise, the customer is assigned to an overflow pool (see more details on feasible overflow assignment in Section 4.1.2).
$s^{1}=s$$s^{\prime}$$\textbf{a}=(a^{1},\dots,a^{q})$$s^{2}$$s^{3}$$s^{q+1}$$\sigma^{1}(s),a^{1}$$\sigma^{2}(s),a^{2}$rrivalDepartureTime $t_{k}$ime $t_{k+1}$
Figure 1: Illustration of the atomic decision process.
Figure 1 illustrates the atomic decision process for deciding and executing the atomic actions. We denote the sequence of customer classes of the chosen order as $\boldsymbol{\sigma}(s)=(\sigma^{1}(s),\sigma^{2}(s),\cdots,\sigma^{q}(s))$, with $\sigma^{n}(s)$ representing the class of the $n$th waiting customer within the chosen order. Starting from the first customer, who is from class $\sigma^{1}(s)$, we execute the atomic action $a^{1}$ for this customer, then we update the state according to $a^{1}$ and sequentially move to the second customer, so on and forth till we execute the atomic action $a^{q}$ for the last waiting customer, who is from class $\sigma^{q}(s)$. The atomic action $a^{n}$ for the $n$th customer is chosen based on the atomic state, $s^{n}$, which is observed after assigning the $(n-1)$th customer. This state has the same structure as the one introduced in Section 3.2, i.e., $s^{n}=(x^{n}_{1},\dots,x^{n}_{J},y_{1},\dots,y_{J},h)$. After executing $a^{n}$, the atomic decision process updates the state to $s^{n+1}$ as
$x^{n+1}_{j}=x^{n}_{j}+\sum_{k\neq j}\mathbbm{1}(a^{n}=j,\sigma^{n}(s)=k)-\sum_{k\neq j}\mathbbm{1}(a^{n}=k,\sigma^{n}(s)=j),\forall j\in\mathcal{J},$ (4)
where $\mathbbm{1}(\cdot)$ is the indicator function. This updated atomic state will be used for deciding $a^{n+1}$.
We note that even though the action is still state-dependent, the largest action space size is $|\mathcal{J}|=J$, which is drastically smaller compared to the combinatorial action space in the original MDP problem. This reduction forms the most critical foundation for developing an algorithm that scales effectively with the number of customer classes and server pools.
For clarity, and to distinguish between the MDP state and action defined in Section 3.2 and those used for the atomic decision process here, we will refer to the state and action in Section 3.2 as the system-level state and action throughout the rest of the paper. Meanwhile, we will refer to the state and action in the atomic decision process as the atomic state and action. It is important to note that the system-level and atomic states and actions are closely related. Specifically, the initial atomic state $s^{1}$ corresponds to the system-level (pre-action) state $s$, and after all atomic actions are executed, the final atomic state $s^{q+1}$ corresponds to the system-level post-action state $s^{+}$. Additionally, the sum of the atomic actions for the inflow and outflow of class $j$ customers can be used to recover the system-level actions. That is, for any given pair $(\boldsymbol{\sigma}(s),\textbf{a})$, we can recover the system-level action $f_{i,j}$ via $\sum_{n=1}^{q}\mathbbm{1}(a^{n}=j,\sigma^{n}(s)=i)$.
4.1.1 Randomized Atomic Action
In this paper, we use randomized policies to prescribe the atomic actions. We will first outline the specifics of this randomization and then discuss its advantages over conventional deterministic policies in Section 4.1.2.
For a given atomic state $s^{n}$, we define a probability distribution over all possible atomic actions for each customer class. We denote this distribution for a class $i$ customer, given $s^{n}$, as $\kappa(\cdot|s^{n},i)$, and the set of distributions as $\kappa=\{\kappa(\cdot|s^{n},i),i\in\mathcal{J}\}$. Consequently, starting from the initial state $s^{1}=s$ (the system-level pre-action state) and following a given order $\boldsymbol{\sigma}(s)$, the probability of selecting a sequence of atomic actions $\textbf{a}=\{a^{1},\dots,a^{q}\}$ is given by $\prod_{n=1}^{q}\kappa(a^{n}|s^{n},\sigma^{n}(s))$, where $(s^{1},\dots,s^{q})$ denotes the corresponding atomic state sequence generated within the atomic decision process, as specified in Equation (4).
According to the correspondence between atomic and system-level actions, the probability of taking a system-level action $f$ under the given distributions $\kappa$ can be expressed as:
$\pi(f|s)=$ $\sum_{\textbf{a}}\prod_{n=1}^{q}\kappa(a^{n}|s^{n},\sigma^{n}(s))\cdot\mathbbm{1}\Big{(}\sum_{n=1}^{q}\mathbbm{1}(a^{n}=j,\sigma^{n}(s)=i)=f_{i,j},\forall i,j\Big{)}.$ (5)
To simplify notation and align with the broader PPO algorithm literature, we use the system-level policy $\pi=\{\pi(f|s)\}$ in the rest of the paper when we refer to a policy. However, it is important to clarify that this system-level policy is driven by the underlying atomic-level probability distributions. In other words, policy $\pi$ is the outcome of the sequence of atomic actions governed by the distributions $\kappa$. Our goal is to optimize $\kappa$ in such a way that the induced system-level policy $\pi$ achieves the performance objective in (3). Moreover, in Section 5.3, we will introduce a “batching” setup that simplifies the calculation of (5) and eliminates the dependence on the specific processing order $\boldsymbol{\sigma}$. This will establish a more direct connection between the system-level policy $\pi$ and the atomic-level distributions $\kappa$. We discuss how the randomized policy can be practically applied in decision-making in Section 6.5.
4.1.2 Policy Parameterization
As discussed, our ultimate goal is to optimize the atomic action probabilities $\kappa$. To achieve this, we parameterize these probabilities using a set of parameters $\theta\in\Theta$. For a given $\theta$, we denote the resulting probabilities for atomic actions across customer classes as $\kappa_{\theta}=\{\kappa_{\theta}(\cdot|s,i),~{}i\in\mathcal{J}\}$. The induced system-level policy, denoted by $\pi_{\theta}$, is then derived from these parameterized atomic probabilities via (5). We refer to $\theta$ as the policy parameter.
While there are various ways to parameterize $\kappa_{\theta}$, in this paper, we use Neural Networks (NN), where $\theta$ represents the NN parameters. Before selecting the atomic action for the $n$th customer, given the atomic state $s^{n}$ and customer class $\sigma^{n}(s)=i$, the NN outputs the probability of choosing $a^{n}=j$ as:
$\kappa_{\theta}\big{(}a^{n}=j|s^{n},\sigma^{n}(s)=i\big{)}=\frac{\text{exp}(g_{i,j}(s^{n}|\theta))}{\sum_{k=1}^{J}\text{exp}(g_{i,k}(s^{n}|\theta))},$ (6)
where $g_{i,j}(s|\theta)$ is one of the $J\times J$ output neurons of a dense NN with input $s$ and parameters $\theta$. In the simplest setting, we can use a NN with no hidden layers, i.e.,
$g_{i,j}\big{(}s|\theta=(w_{i,j},b_{i,j},i,j\in\mathcal{J})\big{)}=w_{i,j}\cdot s+b_{i,j},$
which is equivalent to a multi-class logistic regression model. For further details on NN parameterization, see Section 5.1.
Note that Equation (6) could assign a positive probability to non-feasible atomic actions, such as when the selected server pool is full or cannot serve class $j$. To ensure the feasibility of the prescribed atomic action, we set $\kappa_{\theta}(a^{n}|s^{n},\sigma^{n}(s))=0$ for $a^{n}\not\in\mathcal{A}(s^{n},\sigma^{n}(s))$ (i.e., force the probability of non-feasible actions to 0), and then rescale the remaining probabilities to sum to 1. This adjustment is equivalent to resampling: if a sampled atomic action based on $\kappa_{\theta}$ is not feasible, we continue resampling until a feasible action is selected.
Now, the task of determining an effective overflow policy is reduced to learning the policy parameter $\theta$. We describe how the PPO method can accomplish this goal in the following section. Note that if the NN has $D$ hidden layers and $O(J^{2})$ neurons per layer, the size of the parameter $\theta$ is $O(DJ^{4})$. This is a polynomial growth rate rather than an exponential one, which ensures that the PPO method remains scalable when the number of customer classes and server pools increases.
4.2 Finding Optimal Policy: Proximal Policy Optimization
The randomized policy framework enables the use of policy gradient methods to learn the parameter $\theta$. In this paper, we employ the state-of-the-art policy gradient method, PPO, with custom modifications tailored to our specific queueing network context. We provide an overview of the PPO framework in this section; we defer the details of our tailored modifications to Section 5.
The relative value function under the system-level policy $\pi$ satisfies the Poisson equation
$v_{\pi}(s)=g_{\pi}(s)-\gamma_{\pi}+\mathbb{E}_{s^{\prime}\sim p_{\pi}(s^{\prime}|s)}[v_{\pi}(s^{\prime})],$ (7)
where $\gamma_{\pi}$ denotes the long-run average cost under policy $\pi$, $g_{\pi}(s)=\mathbb{E}_{f\sim\pi(\cdot|s)}g(s,f)$ is the expected one-epoch cost, and $p_{\pi}(s^{\prime}|s)=\mathbb{E}_{f\sim\pi(\cdot|s)}p(s^{\prime}|s,f)$ is the transition probability. Correspondingly, the advantage function of policy $\pi$ is defined as
$A_{\pi}(s,f)=g(s,f)-\gamma_{\pi}+\mathbb{E}_{s^{\prime}\sim p(\cdot|s,f)}[v_{\pi}(s^{\prime})]-v_{\pi}(s),$ (8)
i.e., it measures the relative benefit (cost reduction) of taking a specific action $f$ compared to the expected performance of following the current policy $\pi$ at a given state $s$. In the cost minimization setting, a negative advantage function value means that the action leads to a lower expected cost (which is desirable). Intuitively, policy gradient methods, including PPO, update the policy parameters in the direction that increases the probability of actions with negative advantage values and decreases the probability of those with positive advantage values.
Specifically, assume that the current policy $\pi_{\eta}\in\Pi$ is parameterized by $\eta\in\Theta$, using the atomic action parameterization described in Section 4.1.2. For notational simplicity, let $v_{\eta}$ and $A_{\eta}$ denote the relative value function and advantage function under $\pi_{\eta}$, respectively. To improve upon $\pi_{\eta}$ to get a better policy $\pi_{\theta}$, Schulman et al. (2017) suggests minimizing the following PPO objective:
$L(\theta):=\mathop{\mathbb{E}}\limits_{\begin{subarray}{c}s\sim\mu_{\eta}\\ f\sim\pi_{\eta}(\cdot|s)\end{subarray}}\max\left\{r_{\theta,\eta}(f|s)A_{\eta}(s,f),\quad\text{clip}\left(r_{\theta,\eta}(f|s),1-\epsilon,1+\epsilon\right)A_{\eta}(s,f)\right\}.$ (9)
Here, the distribution $\mu_{\eta}$ denotes the stationary distribution under policy $\pi_{\eta}$; $r_{\theta,\eta}$ denotes the probability ratio between $\pi_{\theta}$ and $\pi_{\eta}$:
$r_{\theta,\eta}(f|s):=\frac{\pi_{\theta}(f|s)}{\pi_{\eta}(f|s)},$ (10)
where $\pi_{\theta}(f|s)$ and $\pi_{\eta}(f|s)$ are defined as in (5), with $\kappa$ replaced by $\kappa_{\theta}$ and $\kappa_{\eta}$, respectively; and the clipped function
$\text{clip}(x,1-\epsilon,1+\epsilon):=\left\{\begin{array}[]{cc}1-\epsilon,&\text{if }x<1-\epsilon\\ 1+\epsilon,&\text{if }x>1+\epsilon\\ x,&\text{otherwise}\end{array}\right..$
The acceptable difference $\epsilon$ is a hyper-parameter that can be tuned for better algorithm performance.
For the theoretical guarantee, we establish Proposition 1, which extends the performance improvement guarantee of PPO from the time-stationary, long-run average cost setting in Dai and Gluzman (2022) to our more complex periodic environment. See the complete statement of Proposition 1 (alongside needed technical conditions) in Appendix 10 and its proof in Appendix 13. The key challenge in the proof lies in the periodicity of state transitions, which makes directly applying previous results inapplicable. To address this, we introduce a new framework using daily Markov chains (MCs), where one period (e.g., one day) is treated as a single transition step. This transformation allows us to redefine the objective function in a way that ensures aperiodicity and enables the application of previous theoretical framework. Proposition 1 establishes that the change in long-run average cost can be decomposed into two terms: one that dominates the improvement and another that diminishes at a faster rate. Since PPO minimizes the dominant term while keeping the policy update conservative to control the secondary term, it leads to a provable policy improvement guarantee.
In implementation, we estimate the objective function (9) using simulation data samples, i.e.,
$\hat{L}(\theta,\mathcal{D}_{\eta}^{T},\hat{\textbf{A}}_{\eta}(\mathcal{D}_{\eta}^{T}))=\frac{1}{Tm}\sum_{t=1}^{T}\sum_{k=0}^{m-1}\max$$\left\{r_{\theta,\eta}(f(t_{k})|s(t_{k}))\hat{A}_{\eta}(s(t_{k}),f(t_{k})),\right.$ (11)
$\left.\text{clip}\left(r_{\theta,\eta}(f(t_{k})|s(t_{k})),1-\epsilon,1+\epsilon\right)\hat{A}_{\eta}(s(t_{k}),f(t_{k}))\right\},$
where the dataset $\mathcal{D}_{\eta}^{T}=\left\{\big{(}s(t_{k}),f(t_{k}),g(t_{k})\big{)}_{k=0}^{m-1}\right\}_{t=1}^{T}$ contains $T$ days of simulation data generated under policy $\pi_{\eta}$, and the vector $\hat{\textbf{A}}_{\eta}(\mathcal{D}_{\eta}^{T})=\left\{\left(\hat{A}_{\eta}(s(t_{k}),f(t_{k}))\right)_{k=0}^{m-1}\right\}_{t=1}^{T}$ consists of the estimated advantage functions for each decision epoch within the $T$ days using the collected data $\mathcal{D}_{\eta}^{T}$. Note that the constant factor $\frac{1}{Tm}$ in $\hat{L}$ can be disregarded without affecting the objective optimization. The detailed steps of the PPO algorithm are given in Algorithm 1.
Algorithm 1 Proximal policy optimization
Output: Policy $\pi_{\theta^{*}}$
Input: Policy $\pi_{\eta}$
1:  for policy iteration $r=0,1,2,...,R-1$  do
2:     Run policy $\pi_{\eta}$ for $T$ days and collect simulation dataset $\mathcal{D}_{\eta}^{T}=\left\{\big{(}s(t_{k}),f(t_{k}),g(t_{k})\big{)}_{k=0}^{m-1}\right\}_{t=1}^{T}$.
3:     Estimate the advantage functions $\hat{\textbf{A}}_{\eta}(\mathcal{D}_{\eta}^{T})$.
4:     Minimize the estimated objective function (11) w.r.t $\theta\in\Theta$ and obtain parameters $\theta^{*}$.
5:     Update $\eta\leftarrow\theta^{*}$.
6:  end for
We conclude this section with two important remarks. First, our policy parameterization differs from the conventional ones used in PPO framework, which would directly parameterize a randomized policy over the entire combinatorial system-level action space. This action space has a size of $\prod_{j=1}^{J}\left(\begin{array}[]{c}q_{j}+w_{j}\\ w_{j}\end{array}\right)$, which means that a conventional policy network would require an extremely large output layer, making training highly challenging. The key element for our algorithm is the combination of atomic actions and the randomized policy. Second, performance improvement using the PPO algorithm is ensured when policy updates are conservative (see Appendix 10). The atomic action probabilities are parameterized using continuous functions that output distributions over small action spaces, allowing incremental updates that are scalable given the relatively small parameter size. This would not be feasible if we use the deterministic policy setup. We have also conducted numerical experiments that demonstrate the superior performance of our randomized approach compared to deterministic ones. The numerical results highlight two major advantages of randomized atomic actions: (i) avoiding chattering – randomized policies integrate well with PPO, which prevent policy “chattering” in discrete-action spaces; (ii) randomized actions naturally allow for exploration, which improve the speed of convergence to a “good” final policy. We discuss how to use the randomized policy in practice in Section 6.5.
5 PPO Design: Tailoring for Periodicity in MDP
A salient feature of our problem setting is the intrinsic time-dependency. The optimal action at different decision epochs within a day can vary significantly, even when the customer counts are similar. While PPO is well-established for stationary environments, applying it to time-nonstationary systems presents additional challenges. Specifically, our problem considers the long-run average cost in a periodic, time-varying setting. This makes ad hoc adjustments such as rolling-horizon methods ineffective. We address the challenges by integrating time-periodicity into the PPO framework and leveraging the queueing system structure. In Section 5.1, we introduce a tailored NN design to represent policies in a periodic setting, capturing both differences and similarities between policies at different epochs. This design is critical for improving sample efficiency (i.e., requiring less simulation data and enabling more efficient computation). In Sections 5.2 and 5.3, we introduce two more elements to further improve algorithm performance and reduce computational time: (i) value function approximations that leverage the queueing dynamics under the randomized atomic policy, and (ii) a batching setup for atomic actions.
5.1 Policy Network Design
As discussed in Section 4.1.1, we use NNs to parameterize the atomic action probabilities $\kappa_{\theta}=\{\kappa_{\theta}(j|s,i),i,j\in\mathcal{J}\}$, which induce the system-level policy $\pi_{\theta}$. We refer to these NNs as the “policy network” and $\theta$ as the NN parameters. To accommodate the periodic setting, we consider three different designs for the policy network, as shown in Figure 2. The first two are intuitive adaptations of policy networks from conventional time-stationary settings, while the third is our own innovation. We describe each design below.
Figure 2: Three policy network structures. The fully-connected network takes epoch index as part of the input and uses a single network to learn actions across all epochs. The fully-separate network trains $m$ separate networks for each of the $m$ epochs. The partially-shared network excludes the epoch index from input but uses different blocks of output neurons for actions of different epochs. The parameters of the first several layers are shared across epochs and while those of the later layers are specific to each epoch.
  • The fully-connected network structure includes the epoch index as part of its input, and generates atomic action probabilities of different epochs with the same set of NN parameters. Specifically, given state $s\in\mathcal{S}$, we use one-hot encoding to include the epoch index $h$ in the input, encoding $s$ as $(\tilde{s},e^{h})$, where $\tilde{s}=(x_{1},\dots,x_{J},y_{1},\dots,y_{J})$ is the state excluding epoch index, and $e^{h}\in\mathbb{R}^{m}$ is the unit vector with the $h$th element being 1 and all others as 0. The action probabilities are then given by:
    $\kappa_{\theta}(j|s=(\tilde{s},h),i)=\frac{\text{exp}(g_{i,j}(\tilde{s},e_{h}|\theta))}{\sum_{k=1}^{J}\text{exp}(g_{i,k}(\tilde{s},e_{h}|\theta))},\quad i,j\in\mathcal{J},s\in\mathcal{S},$
    where $g_{i,j}$ is one of the $J\times J$ output neurons of the fully-connected network. This design is the most straightforward way to handle time-dependency – the network uses the same set of parameters to generate outputs for different epochs, which captures temporal differences by incorporating time as part of the input.
  • The fully-separate network structure uses $m$ independent networks for each of the $m$ different epochs, so the epoch index does not need to be part of the input. Specifically, the NN parameters $\theta=(\theta^{0},\dots,\theta^{m-1})$, where $\theta^{h}$ is the parameters for the network corresponding to epoch $h$. The atomic action probabilities are given by:
    $\kappa_{\theta}(j|s=(\tilde{s},h),i)=\frac{\text{exp}(g_{i,j}^{h}(\tilde{s}|\theta^{h}))}{\sum_{k=1}^{J}\text{exp}(g_{i,k}^{h}(\tilde{s}|\theta^{h}))},\quad i,j\in\mathcal{J},s\in\mathcal{S},$
    where $g_{i,j}^{h}$ is one of the $J\times J$ output neurons of the network for epoch $h$. This structure allows each epoch to have its own independent set of network parameters to capture the differences between them, but it does not exploit any potential similarities across epochs.
  • The partially-shared network structure is a new structure we propose in this paper. This structure takes a middle ground between the fully-connected and fully-separate structures: it excludes the epoch index from the input while structuring the NN parameters such that the first several layers are shared across epochs and the later layers are specific to each epoch. Specifically, we define the NN parameters as $\theta=(\tilde{\theta},\theta^{1},\dots,\theta^{m})$, where $\tilde{\theta}$ represents the shared parameters (typically for the first few layers), and $(\theta^{1},\dots,\theta^{m})$ correspond to the epoch-specific parameters (typically for the final layer). The atomic action probabilities are given by:
    $\kappa_{\theta}(j|s=(\tilde{s},h),i)=\frac{\text{exp}(g_{i,j}^{h}(\tilde{s}|\tilde{\theta},\theta^{h}))}{\sum_{k=1}^{J}\text{exp}(g_{i,k}^{h}(\tilde{s}|\tilde{\theta},\theta^{h}))},\quad i,j\in\mathcal{J},s\in\mathcal{S}.$
    In other words, this structure does not differentiate states across epochs in the input but selects the corresponding block of output neurons for each epoch. This allows us to leverage data from all epochs to train the network while still prescribing different action probabilities for different times of the day.
To ensure feasibility, given state $s=(\tilde{s},h)$, for a non-feasible pool $k\not\in\mathcal{A}(s,j)$ of class $j$, we can set $g_{j,k}(s|\theta)=0$ for the fully-connected structure, $g_{j,k}^{h}(\tilde{s}|\theta^{h})=0$ for fully-separate, and $g_{j,k}^{h}(\tilde{s}|\tilde{\theta},\theta^{h})=0$ for partially-shared, followed by re-normalizing the remaining probabilities.
Combining advantages. The fully-connected and fully-separate designs each have their pros and cons, while the partially-shared policy network combines their strengths. Specifically, the fully-connected network design uses the same set of parameters across all time epochs, naturally capturing similarities between epochs and enabling training with data from all time epochs. However, since the policy differences between epochs can be highly complex, a more sophisticated network (e.g., with additional hidden layers) may be needed to capture these nuances and hence more training data. The fully-separate network design accounts for differences between epochs with independent networks, but it would also require more time and data to train, as each network is only trained using data from its corresponding epoch. The partially-shared network design strikes a balance: shared parameters in the early layers capture commonalities between epochs for efficient training, while the separated last layer adapts to epoch-specific policy outputs. In the case study of a ten-pool system, the partially-shared design requires only a few thousand simulation data samples for training, yet outperforms the other two designs that need 100 times more data. We will provide a detailed comparison of the sample efficiency of these three designs in Section 6.4.
5.2 Policy Evaluation: Linear Value Function Approximation
In the standard PPO implementation, a separate neural network, known as the value network, is used alongside the policy network to approximate value functions. The value network is trained with targets estimated via Monte Carlo evaluation. These estimates can have high variance in long-run average settings, which require large amounts of data to achieve accurate policy evaluation. To address this, we adopt a linear value function approximation and leverage the queueing system structure to design proper basis functions. This approach mitigates variance-related inefficiencies and contrasts with the black-box nature of NN-based value estimation. Additionally, our approach extends to other operations problems with similar structural properties.
The key to accurate linear value function approximations lies in the choice of basis functions. Our proposed basis functions are:
$V_{d}(s),~{}x_{j},~{}x_{j}^{2},~{}y_{j},~{}y_{j}^{2},~{}x_{j}y_{j},\quad j\in\mathcal{J}.$ (12)
Here, the polynomial terms are the same as those in Dai and Shi (2019). We introduce a new design, $V_{d}(s)=\sum_{j=1}^{J}\hat{v}_{\pi,j}(s_{j})$, which approximates the relative value function under the randomized atomic policy. Each $\hat{v}_{\pi,j}(s_{j})$ relies only on pool-$j$’s state information, $s_{j}=(x_{j},y_{j},h)$, and estimates the relative value function based on the holding and overflow cost components for class-$j$ customers.
Specifically, under the randomized atomic policy $\pi$, we decompose the original $J$-pool system into $J$ single-pool systems. For each decomposed pool-$j$ system, the relative value function $\hat{v}_{\pi,j}$ is solved via the corresponding post-action Poisson equation
$\hat{v}_{\pi,j}(s_{j})=C_{j}q_{j}-\hat{\gamma}_{\pi,j}+\mathbb{E}\left[\sum_{k=1}^{J}B_{j,k}\bar{f}_{j,k}+\hat{v}_{\pi,j}(s_{j}^{\prime})\right]$ (13)
for a given post-action state $s_{j}$ (we drop the superscript $+$ here for simplicity). Here, $C_{j}q_{j}$ is the post-action holding cost with $q_{j}=\max(s_{j}-N_{j},0)$, $\hat{\gamma}_{\pi,j}$ is the average cost of the decomposed pool-$j$ system, and the cost-to-go term consists of: (a) the expected overflow cost $\mathbb{E}[\sum_{k=1}^{J}B_{j,k}\bar{f}_{j,k}]$, where $\bar{f}_{j,k}$ approximates the overflow assignments from class $j$ to other pools, and (b) the expected value function at the next post-action state $\mathbb{E}\left[\hat{v}_{\pi,j}(s’_{j})\right]$. To evaluate (a) and (b), we need to properly adjust the queueing dynamics for each decomposed pool-$j$ system by incorporating overflow assignments from other classes while excluding overflow assignments from class $j$ to other pools (treated as “diversions”). To achieve this, we approximate the probability of assigning a class-$j$ customer to pool $k$ under policy $\pi$ using $\bar{\kappa}_{j}(k|s_{j},j)$, which depends only on pool-$j$’s state information. Given the arrivals are Poisson and leveraging the Poisson thinning property, the outflows $\bar{f}_{j,k}$ and inflows $\bar{f}_{i,j}$ can be approximated as Poisson distributions with rates $\lambda_{j}\bar{\kappa}_{j}(k|s_{j},j)$ and $\lambda_{i}\bar{\kappa}_{i}(j|s_{j},i)$, respectively. This formulation allows efficient evaluation of both (a) and (b). Details on computing $\bar{\kappa}_{j}$, deriving $\hat{v}_{\pi,j}$, and formulating (13) using post-action states are provided in Appendix 11.
The feasibility of the pool-wise decomposition relies on the randomized atomic policy. The continuous nature of overflow probabilities under this policy allows us to generate good approximations of the $f_{i,j}$ values. In contrast, a deterministic policy results in discrete actions with more drastic variations across states, making it harder to approximate overflow proportions effectively. This distinction makes our new basis function, $V_{d}$, specifically suited for the PPO framework and different from the $V_{s}$ basis function proposed in Dai and Shi (2019), which was designed for deterministic, discrete-action spaces. In Section 6.4, we compare the performance of $V_{d}$ and $V_{s}$ alongside other polynomial bases. Our results demonstrate that $V_{d}$ improves sample efficiency and value function approximation accuracy, which ultimately leads to better policy performance.
5.3 Batching Setup and Other Implementation Details
In the general atomic decision process, each of the atomic actions sequentially depends on the atomic states as illustrated in Figure 1. We propose a batching setup to simplify this process, which uses the same initial state $s=s^{1}$ for all atomic actions in a batch. Specifically, given $s$ and a chosen customer class order $\boldsymbol{\sigma}(s)$, the probability of conducting atomic action sequence a becomes $\prod_{n=1}^{q}\kappa(a^{n}|s,\sigma^{n}(s))$. In other words, during this batch, since the input state $s$ no longer changes, we prescribe atomic actions to the customers from the same class $i$ following the same distribution $\kappa(\cdot|s,i)$. Under the assumption that any system-level action $f$ satisfying $\sum_{j}f_{i,j}=q_{i},\forall i$ is feasible, we have the following lemma.
Lemma 5.1
Under the batching setup, for a given system-level state $s$ and any pre-determined rule for choosing customer order $\boldsymbol{\sigma}(s)$, the probability of taking action $f$ follows a multinomial distribution that is independent of the order $\boldsymbol{\sigma}(s)$, i.e.,
$\pi(f|s)=\prod_{i=1}^{J}\frac{q_{i}!}{\prod_{j=1}^{J}f_{i,j}!}\prod_{j=1}^{J}\kappa(j|s,i)^{f_{i,j}}.$ (14)
Proof 5.2
Proof. From (5), we have that
$\pi(f|s)=$ $\sum_{\textbf{a}}\prod_{n=1}^{q}\kappa(a^{n}|s,\sigma^{n}(s))\cdot\mathbbm{1}\left(f_{i,j}=\sum_{n=1}^{q}\mathbbm{1}(a^{n}=j,\sigma^{n}(s)=i),i,j\in\mathcal{J}\right)$
$=$ $\left(\prod_{i=1}^{J}\prod_{j=1}^{J}\kappa(j|s,i)^{f_{i,j}}\right)\cdot\sum_{\textbf{a}}\mathbbm{1}\left(f_{i,j}=\sum_{n=1}^{q}\mathbbm{1}(a^{n}=j,\sigma^{n}(s)=i),i,j\in\mathcal{J}\right)$
$=$ $\prod_{i=1}^{J}\frac{q_{i}!}{\prod_{j=1}^{J}f_{i,j}!}\prod_{j=1}^{J}\kappa(j|s,i)^{f_{i,j}}.$
Here, the second equality uses the fact that the atomic actions for class $i$ follow the same distribution $\kappa_{\theta}(\cdot|s,i)$, so that we can group the atomic actions for each of the $q$ waiting customers simply by their class $i$ and the assigned pool $j$, which connects to the $f_{i,j}$. The last equality comes from the combinatorial partition of assigning the $q_{i}$ waiting customers from class $i$ to each pool $j$. $\square$
This batching setup is new in our application context compared to other atomic action approaches (e.g., Feng et al. (2021)), as our system-level overflow decisions are made on a discrete-time basis with multiple customers to be processed in a batch. The first advantage of this batching approach is that the atomic policy for the same class of customers remains consistent within the same decision epoch, which ensures a degree of fairness regardless of the pre-selected processing order $\boldsymbol{\sigma}(s)$. The second advantage is the improved time efficiency in both the simulation and policy update processes. Unlike the original sequential atomic decision setup, we do not need to update each atomic state or compute the state-dependent atomic action probabilities for each individual customer. The third advantage lies in the great simplification of calculating the objective function (11) for PPO. Specifically, recall that we need to compute the probability ratio
$r_{\theta,\eta}(f|s)=\frac{\pi_{\theta}(f|s)}{\pi_{\eta}(f|s)}$
at each decision epoch, where $\pi_{\theta}(f|s)$ and $\pi_{\eta}(f|s)$ are defined as in (5), with $\kappa$ replaced by $\kappa_{\theta}$ and $\kappa_{\eta}$, respectively. Under the original setup, getting this ratio is computationally intensive because it involves summing over all possible atomic action sequences $\boldsymbol{a}$ corresponding to the system-level action $f$ for a given customer order $\boldsymbol{\sigma}(s)$ (i.e., the summation over $\boldsymbol{a}$ part in (5)). In contrast, under the batching setup, using (14) from Lemma 14, this ratio is greatly simplified to:
$r_{\theta,\eta}(f|s)=\prod_{i=1}^{J}\prod_{j=1}^{J}\left(\frac{\kappa_{\theta}(j|s,i)}{\kappa_{\eta}(j|s,i)}\right)^{f_{i,j}},$
since the combinatorial factor $\frac{q_{i}!}{\prod_{j=1}^{J}f_{i,j}!}$ is a constant for given $f_{i,j}$’s and cancels out in the ratio calculation. This also establishes a direct equivalence between the ratio of atomic action probabilities ($\kappa_{\theta}$, $\kappa_{\eta}$) and the ratio of the induced system-level policies ($\pi_{\theta}$, $\pi_{\eta}$).
As in the original setup, it is possible that during implementation a sampled atomic action under the batching setup could become infeasible if the selected pool no longer has an available server or cannot accommodate the customer class. To handle this, we apply the same resampling technique to ensure a feasible action is selected. Other implementation details are as follows. We use Tensorflow v1 to build the policy network training pipeline and use the Ray package for parallel simulation data generation. The PPO algorithm iterates through three major modules: simulation data generation, policy evaluation, and policy improvement. For data generation, multiple actors (CPUs) run parallel simulations of equal length, and then the simulated data is concatenated into one long episode. For policy evaluation, we compute the linear value function approximations and the estimated advantage functions. For policy improvement, we update the policy network parameters $\theta$ using the Adam optimization method, running for $E$ training epochs (one training epoch corresponds to the number of complete passes through the entire training dataset). This process repeats until the average costs from two iterations differ by less than a small $\delta$. Table 1 summarizes important PPO hyper-parameters used in a ten-pool system in the case study in Section 6. More details on tuning these hyper-parameters are provided in Appendix 12.1.
Parameters Baseline Choice
Neural network structure Partially-shared
Neural network complexity One hidden layer with $34$ neurons
Basis function $(V_{d},X,X^{2},Y,Y^{2},XY)$
Initial policy Complete-overflow
Simulation days per actor 10,000
Number of actors 10
Number of training epochs $(E)$ 15
Clipping parameter $(\epsilon)$ 0.5
Gap tolerance $(\delta)$ 0.1
Table 1: Baseline choice of PPO hyper-parameters in the ten-pool system.
6 Case Study
In this section, we evaluate the effectiveness of our tailored PPO algorithm through a case study on inpatient overflow assignment in large hospital systems, including settings with up to twenty patient classes and twenty wards (pools). In Section 6.1, we introduce the case study settings. In Section 6.2, we show that our PPO algorithm achieves similar performance to the ADP method from Dai and Shi (2019) – considered the “state-of-the-art” – while significantly reducing computational time in the five-pool baseline, the largest system handled by Dai and Shi (2019). In Section 6.3, we focus the analysis on ten- and twenty-pool systems, where our method significantly outperforms several benchmark policies. In Section 6.4, we examine the algorithm’s performance across different hyper-parameter settings and highlight the critical role of our atomic policy design and tailored components. Finally, in Section 6.5 we summarize key takeaways and offers practical recommendations for managing inpatient overflow.
6.1 Model Settings
The model settings for our case study are based on the five-pool system from Dai and Shi (2019), which is calibrated using aggregate patient-flow statistics from a large teaching hospital in Asia. The calibrated system has 315 servers (inpatient beds) and a daily arrival rate of 70 customers (patients). Time-varying arrival and discharge patterns are detailed in Appendix 12. Each class (medical specialty) has a dedicated primary ward, one preferred overflow ward, and two secondary overflow wards. The corresponding overflow costs are $B=(B_{1},B_{2})$, with $B_{1}<B_{2}$ to reflect ward preferences. We set the baseline values at $B=(30,35)$ for overflow costs and $C=6$ for holding costs. These cost parameters act as “tuning knobs,” which allow managers to adjust them to generate Pareto curves and balance congestion with overflow rates to meet target performance goals.
We extend this model to ten-pool and twenty-pool systems. The ten-pool system consists of two departments: a VIP department with smaller capacity and a regular department with larger capacity. Both departments follow the five-pool structure, and the total capacity is double that of the baseline five-pool system. Arrival rates are adjusted proportionally, as shown in Table 2. Overflow routes within each department remain unchanged (one preferred and two secondary overflow wards), but we also allow VIP patients to overflow to the regular department (not vice versa). Figure 3 provides a schematic representation of the feasible routes in the ten-pool system, using the GeMed (General Medicine) specialty as an example. The diagram illustrates both within-department routes (black solid lines and red/blue dashed lines) and cross-department routes (orange and green dotted lines), highlighting the overflow pathways available within and between departments. Overflow costs are set at $B=(B_{1},B_{2},B_{3},B_{4})=(25,30,35,40)$, with lower costs for within-department overflows $(B_{1},B_{2})$ and higher costs for cross-department overflows $(B_{3},B_{4})$. The unit holding cost is set at $C=6$ for regular patients and $C=7$ for VIP patients. For the twenty-pool system, we assume it consists of two hospitals, each modeled as a ten-pool system with different capacities, as shown in Table 2. Patients can overflow within their hospital or to the other hospital. Overflow costs are set at $B=(B_{1},B_{2},B_{3},B_{4},B_{5},B_{6},B_{7},B_{8})=(25,30,35,49,40,45,50,55)$, with lower costs for within-hospital overflows $(B_{1},\dots,B_{4})$ and higher costs for cross-hospital overflows $(B_{5},\dots,B_{8})$. The baseline unit holding costs remain the same, i.e., $C=6$ for regular and $C=7$ for VIP patients.
[…]
Figure 3: Illustration of the ten-class ten-pool system based on practice in our partner hospital. The solid black arrow depicts the primary assignment route, while red and blue dashed arrows show preferred and secondary overflow assignments within a department. Additionally, orange and green dotted arrows represent cross-department overflow assignments (only from VIP to regular).
Number of inpatient beds Daily arrival rate Nominal utilization
Five-pool 315 70 0.889
Ten-pool VIP dept. 210 50 0.962
Regular dept. 420 90 0.857
Total 630 140 0.889
Twenty-pool Hospital 1 VIP dept. 175 50 1.143
Regular dept. 385 90 0.835
Total 560 140 1
Hospital 2 VIP dept. 245 50 0.816
Regular dept. 455 90 0.791
Total 700 140 0.800
Total - 1260 280 0.889
Table 2: Parameter settings for five-pool, ten-pool and twenty-pool models.
6.2 Five-pool System: Comparing Performance with ADP
Table 3 compares the performance of our PPO algorithm with the ADP method in the five-pool model. We consider two setups: a balanced setting where each pool has equal capacity, and an unbalanced setting where some wards have less capacity and higher congestion. Each day is divided into eight decision epochs, and following Dai and Shi (2019), we fix the unit holding cost while varying the unit overflow cost. These experiments show that the PPO method achieves performance comparable to ADP (within 4% difference) – the best-performing method for the five-pool model to date – while significantly reducing computational time. Specifically, PPO completes one iteration in about two hours, compared to over ten hours for ADP, since its atomic policy design eliminates the need for exhaustive search over all feasible actions. Furthermore, in the simpler two-pool systems where the exact optimal policy is known, PPO achieves near-optimal results (within 3% difference). More results are available in Appendix 12.
Eight-epoch; Unbalanced Eight-epoch; Balanced
B=(15,25) B=(30,35) B=(40,45) B=(15,25) B=(30,35) B=(40,45)
ADP 152.68$\pm$0.76 196.31$\pm$0.92 250.39$\pm$0.45 132.46$\pm$0.32 193.03$\pm$0.62 248.48$\pm$0.71
PPO 153.17$\pm$0.34 202.37$\pm$0.65 260.87$\pm$0.40 134.59$\pm$0.82 196.78$\pm$1.06 258.32$\pm$0.98
Table 3: Five-pool system. We set $N_{1}=60,N_{2}=64,N_{3}=67,N_{4}=N_{5}=62$ in unbalanced setting and $N_{1}=\cdots=N_{5}=63$ in balanced setting. Original atomic decision (non-batched) is used here.
6.3 Large Systems: Algorithm Scalability
In larger systems with more than five pools, the ADP method becomes infeasible due to the need to exhaustively search over the combinatorially large action space. Therefore, we compare the performance of the PPO algorithm against three benchmark policies introduced in Dai and Shi (2019): (a) the complete-overflow policy, where overflow assignments are made as many as possible at each decision epoch; (b) the midnight policy, allowing overflow assignments only at midnight each day; and (c) the empirical policy, which reflects current practice by allowing overflow assignments only during nighttime (7 PM to 7 AM the next day). Figure 4 shows the performance comparison in ten-pool and twenty-pool systems, where PPO significantly outperforms all three naive policies. In the ten-pool system, PPO achieves an average cost of 300, approximately 23% lower than the best-performing benchmark (the empirical policy with an average cost of 390). For the twenty-pool system, the improvement is similar: PPO reduces the average cost to 1000, about 25% lower than the empirical policy (which has an average cost of 1250). The computational time is about 3.5 hours per iteration for the ten-pool system and about 8 hours per iteration for the twenty-pool systems. Note these are training times - once the policy networks are trained, the execution of the optimal policy is almost instantaneous in real-time deployment. These results highlight the scalability and effectiveness of our PPO algorithm in large-scale systems.
(a) Ten-pool
(b) Twenty-pool
Figure 4: Performance comparison in the ten-pool and twenty-pool systems.
6.4 Importance of Tailored Designs
In this section, we evaluate the performance of different hyper-parameter combinations in the ten-pool system, using the values from Table 1 as the baseline. While hyper-parameters for NNs play a role in algorithm performance as traditionally understood, we find that they are less critical than the design of the policy network, basis functions, and batching setup – the tailored elements introduced in this paper. These results highlight the importance of adapting a general-purpose algorithm to the specific problem domain for optimal performance and computational efficiency.
Policy network design. We compare the performance of three policy NN designs in the ten-pool system across various neural network widths, depths, and training settings; see Figures 5-6. The figures plot the percentage improvement in average cost achieved by the PPO policy (after 10 iterations) compared to the best-performing benchmark, the empirical policy, thus, a higher percentage indicates greater improvement (more desirable). We make the following observations.
First, the partially-shared network, represented by the solid (orange) line, consistently outperforms the other two designs across all settings. This holds even when the partially-shared network is trained with (i) fewer simulation days (1,000-5,000 days) compared to $\geq 50,000$ days for the other designs, (ii) fewer training epochs (5 epochs) compared to $\geq 10$ epochs for the other designs, and (iii) simpler NN structures (1 hidden layer with 17 neurons) versus more complex networks. These findings emphasize that the choice of policy network representation has a first-order effect—it significantly impacts algorithm performance improvement. Moreover, the results suggest that our partially-shared design effectively reduces average costs while also cutting down computational time, as it is more sample-efficient and requires a less complex NN to train.
Second, Figure 5 shows that the fully-separate design is the most sensitive to increases in simulation days and training epochs among the three designs. It is because this design learns a separate policy for each epoch and requires significantly more data and longer training to achieve good performance. In contrast, Figure 6 indicates that the fully-connected design benefits from increased network depth, since deeper networks better capture the complex differences across epochs while using a shared policy representation. These numerical findings align with the advantages and limitations of each design discussed in Section 5.1.
(a) Impact of simulation days
(b) Impact of training epochsWhen training data is very small (1,000 simulation days per actor), to achieve good performance within fewer iterations, we use the no-overflow policy as the initial policy (instead of complete-sharing in the baseline) and set the training epochs to 30 (instead of 15 in the baseline).
Figure 5: Trade-off between performance and training days and epochs. Performance is measured as the % improvement in average cost between the PPO policy and the best-performing benchmark (the empirical policy).
(a) Impact of NN depth
(b) Impact of NN width
Figure 6: Trade-off between performance and NNs’ depth and width. Performance is measured as the % improvement in average cost between the PPO policy and the best-performing benchmark (the empirical policy). A value of “0” in the depth or width indicates no hidden layer, which is possible for fully-connected or fully-separate networks but not for the partially-shared network due to its design structure.
Other tailored designs. Table 4 compares the average costs from using our proposed basis function against those used in Dai and Shi (2019), with varying amounts of simulation data. The table shows that our proposed basis functions are much less sensitive to the number of simulation days per actor. While the results with a larger dataset (10,000 days) are similar between the two sets of basis functions, the advantage of our proposed one becomes more evident with fewer simulation days (1,000 days) – the more practical choice for real-world implementation of the algorithm. Moreover, the results in Appendix 12.2 show that our proposed basis functions consistently obtains more accurate approximation among the top 25% most frequently visited states.
Batching is another key design for improving time efficiency, as discussed in Section 5.3. We implement the original atomic decision process following Feng et al. (2021) for our problem and compare it with our batching setup. The batching setup delivers similar performance to the original but reduces computational time by approximately 50% (e.g., from two hours to one hour); see Appendix 12.3. These findings show the value of our tailored designs in enhancing the time efficiency of the PPO algorithm. That is, by adapting to the queueing structure under the randomized atomic action setup, these tailored designs reduce sensitivity to the amount of simulation data. This not only speeds up the algorithm significantly but also makes it more practical for real-world applications where large-scale simulation data may be limited.
Basis function 1k days per actor 10k days per actor
Proposed 311.75$\pm$3.74 302.01$\pm$1.28
Dai and Shi (2019) 325.71$\pm$4.07 307.71$\pm$1.31
Table 4: Comparing the average costs from PPO algorithm using different basis functions.
Impact of other parameters. We present in Appendix 12.1 a detailed performance comparison under various combinations of hyper-parameters. The key takeaway remains consistent with our earlier discussion: our tailored PPO design captures the most salient features of the problem and is of first-order importance in reducing long-run average costs and improving computational efficiency. While tuning other hyper-parameters, such as the number of simulation days, training epochs, and policy update iterations, can further enhance performance, their impact is relatively marginal.
From the comparison results, we recommend increasing training epochs before expanding simulation data or iterations to maximize the use of existing data. Additionally, we identify two key techniques that further improve the algorithm’s efficiency:
  1. 1.
    Reuse of simulation data: Since PPO allows only small policy changes in each iteration, simulation data from a few previous iterations can be reused to reduce variance in estimating the value functions in the current iteration without introducing significant bias. This improves sample efficiency. Our experiments suggest that reusing data from the previous three iterations achieve the best results; see Table 12 in Appendix 12.3.
  2. 2.
    Adaptive clip size: A larger clip size can speed up policy improvement by allowing more substantial changes early in the training. However, it may hinder convergence to a stable policy later. Our experiments recommend starting with a larger clip size (e.g., 0.5) and reducing it as the policy update progresses; see Table 13 in Appendix 12.3.
6.5 Practical Implementation
A key strength of our approach is its domain adaptation: leveraging queueing structures and problem-specific characteristics to develop more effective and efficient algorithms. By integrating a tailored policy network, customized basis functions, and a batching setup, our PPO algorithm requires significantly less simulation data than typical reinforcement learning or ADP-based approaches. This not only accelerates training but also makes the algorithm more practical in environments with limited data or computational resources. The most important insights from our case study findings lies here as well, that is, the domain-aware adaptation is more critical to algorithm performance than fine-tuning neural network parameters in the original general-purposed version, which can be ad hoc and lack systematic guidance.
For practical implementation, since our algorithm is model-based, the hospital first needs to calibrate the queueing system to reflect specific patient flows, arrival rates, and discharge patterns. Such calibration has been successfully applied in various hospital settings, e.g, Bertsimas and Pauphilet (2024). Once calibrated, managers can vary parameters such as holding and overflow costs, which act as levers to balance competing objectives: minimizing patient waiting or reducing overflow occurrences. By adjusting these values, managers can explore trade-offs and generate Pareto-optimal solutions to meet their desired operational goals. For instance, increasing overflow costs could reduce the amount of overflow, or vice versa.
Note that the PPO algorithm employs a randomized policy that provides action probabilities for patient routing decisions. Decision-makers can use these probabilities, $\pi(f|s)$, to guide overflow strategies, such as selecting the most likely action or using de-randomization techniques (e.g., adding a final layer to the policy network for deterministic choices). The randomized policy also offers flexibility for managers to adjust routing decisions in real time. For example, while the algorithm suggests the broad direction such as which wards or departments require the most overflow, managers can refine these suggestions based on real-world factors like patient conditions or staff availability. Ultimately, decisions regarding which individual patients to overflow will depend on the judgment of hospital managers, hospitalists, and charge nurses; our algorithm supports, rather than replaces, the overflow decision-making.
7 Conclusion
This paper presents a scalable and efficient PPO algorithm for managing patient overflow in large hospital systems. Overflow management becomes increasingly complex as the number of hospital wards and patient classes increases, imposing challenges that traditional methods like ADP struggle to handle at scale. Our approach addresses this scalability issue by leveraging atomic actions and a partially-shared policy network. Additionally, our tailored designs for the hospital-specific queueing structures allow the training of our algorithm efficient and reduce the need for extensive simulation data, which overcomes a common limitation of reinforcement learning algorithms. The case study results demonstrate that our PPO algorithm matches or outperforms existing benchmarks in systems with up to twenty hospital wards.
There are several promising directions for future research building on the framework developed in this paper. One potential area is exploring the integration of patient-level contextual data, such as individual health conditions or treatment plans, to further personalize overflow decisions for better patient outcomes. Another avenue of interest is incorporating uncertainty in patient arrival rates, e.g., through online learning, to adapt to sudden surges in demand during public health crises. Moreover, our algorithm testing has thus far been on simulated systems, an essential first step in algorithm development. We are actively seeking hospital partners to deploy the algorithm in real-world settings and gather pilot data. Incorporating a human-in-the-loop component will be crucial for balancing algorithmic recommendations with human decision-making to make the decision-support adaptable to the complexities of hospital operations.
[…]

### v4

@@
 […]
-$0$
+0000
 […]
@@
 Current practices rely on ad-hoc rules, while prior analytical approaches struggle with the intractably large state and action spaces inherent in patient-unit matching.
-A scalable decision-support framework is needed to optimize overflow management while accounting for time-periodic fluctuations in patient flow.
+Scalable decision support is needed to optimize overflow management while accounting for time-periodic fluctuations in patient flow.
 Methodology/Results: We develop a scalable decision-making framework using Proximal Policy Optimization (PPO) to optimize overflow decisions in a time-periodic, long-run average cost setting.
-To address the combinatorial complexity, we introduce atomic actions, which decompose multi-patient routing into sequential, tractable assignments.
+To address the combinatorial complexity, we introduce atomic actions, which decompose multi-patient routing into sequential assignments.
 We further enhance computational efficiency through a partially-shared policy network designed to balance parameter sharing with time-specific policy adaptations, and a queueing-informed value function approximation to improve policy evaluation.
@@
 Case studies on hospital systems with up to twenty patient classes and twenty wards demonstrate that our approach matches or outperforms existing benchmarks, including approximate dynamic programming, which is computationally infeasible beyond five wards.
-Managerial Implications: Our framework offers a scalable, efficient, and adaptive solution for managing patient flow in complex hospital systems.
-More broadly, our results highlight that domain-aware adaptation – leveraging queueing structures and operational insights – is more critical to improving algorithm performance than fine-tuning neural network parameters when applying general-purpose algorithms to specific domain applications.
+Managerial Implications: Our framework offers a scalable, efficient, and explainable solution for managing patient flow in complex hospital systems.
+More broadly, our results highlight that domain-aware adaptation is more critical to improving algorithm performance than fine-tuning neural network parameters when applying general-purpose algorithms to specific applications.
 […]
-[…]
+Time-Periodic Multi-class Queue; Patient Flow; Scalable Decision-Making
 1 Introduction
@@
 1.1 Paper Overview and Contribution
-This paper addresses the existing gap by developing an algorithmic framework based on the Proximal Policy Optimization (PPO) method to minimize the long-run average costs associated with overflow.
-Our key contributions are summarized as follows.
+This paper addresses the existing gap by developing an algorithmic framework based on the Proximal Policy Optimization (PPO) method. Our key contributions are summarized as follows.
   1. 1.
@@
 Beyond healthcare, our framework is applicable to a broad class of matching problems in operations management with similar queueing structures and contributes to efficiently managing the combinatorial complexity in customer-server matching across domains.
-    The success of this framework hinges on two key elements: (a) leveraging the “atomizable” action space and (b) incorporating domain knowledge of queueing systems – particularly in designing value function approximation and policy representation.
-We elaborate on them next.
+    The success of this framework hinges on two key elements: (a) leveraging the “atomizable” action space and (b) incorporating domain knowledge of queueing systems.
   2. 2.
-    Action Atomization: We propose the atomic action in Section 4.1, which decomposes class-level decisions into sequential individual-level assignments.
-This transformation fundamentally simplifies the original combinatorial action space and allows for efficient policy optimization via gradient-based methods.
-Despite its seemingly myopic nature, we demonstrate in our case study that this approach achieves performance comparable to brute-force optimization over the entire action space (Dai and Shi 2019), i.e., this approach works.
-Moreover, it retains flexibility for aggregate planning and incorporating individual preferences for practical implementation.
+    Action Atomization: As detailed in Section 4.1, atomic action decomposes class-level decisions into sequential individual-level assignments, which is critical for simplifying the original combinatorial action space.
+Despite its seemingly myopic nature, we show that this approach, when combined with PPO, achieves performance comparable to the ADP benchmark in a five-pool system, the largest setting considered in Dai and Shi (2019); i.e., this approach works.
+The atomic formulation is inspired by Feng et al.
+(2021), who coined the term atomic action decomposition in the context of finite-horizon ride-hailing problems.
+Extending this idea to the more challenging infinite-horizon, long-run average cost setting is nontrivial, and to the best of our knowledge, this paper is the first to show that combining atomic actions with PPO is effective in this regime.
+We discuss related developments in this growing research area in Section 2.
   3. 3.
-    Dealing with Time-periodicity: To account for real-world daily and weekly periodic patterns in patient arrivals and discharges, we integrate time-periodicity into our PPO framework by leveraging the queueing system structure.
+    Dealing with Time-periodicity: To account for real-world daily and weekly periodic patterns in patient arrivals and discharges, we integrate time-periodicity into our PPO framework.
 First, we establish a theoretical guarantee for policy improvement of PPO under the periodic setting (Section 4.2).
 Second, we design a partially-shared policy network that balances parameter sharing and temporal specificity (Section 5.1).
-This design captures structural similarities while being adaptive to time variations.
-It outperforms two conventional alternatives: (i) a fully-shared network that requires more complex policy networks to model time dependencies, and (ii) a fully-separate network that demands significantly more training data.
-Lastly, we propose a new value function approximation derived from the Poisson equation and tailored to the queueing dynamics under the atomic actions – it decomposes the system-wide value function into pool-wise approximations to improve computational efficiency and stability (Section 5.2).
-Case study results in Section 6 confirm that our framework, which integrates all these design elements, enhances sample efficiency while maintaining superior performance over other benchmarks.
+Lastly, we propose a new value function approximation tailored to the queueing dynamics under the atomic actions to improve computational efficiency (Section 5.2).
+  4. 4.
+    Actionable Insights: Case study results in Section 6, including ablation studies, confirm that our integrated framework enhances sample efficiency while maintaining superior performance over other benchmarks.
+These results highlight that domain-aware algorithm design is more impactful than tuning generic hyperparameters.
+Moreover, our policy visualization and interpretability analysis reveal that the learned policies are both intuitive (aligning with managerial expectations) and insightful (uncovering nuanced network effects that are difficult to capture with heuristic rules).
+These insights provide actionable guidelines for managing overflow while also highlight the necessity of sophisticated, data-driven tools in complex hospital networks.
 2 Literature review
@@
 Adding new beds to alleviate inpatient congestion is often not feasible due to constraints related to space, finances, and regulations.
-Therefore, we focus on literature that explores strategies for utilizing existing bed resources more efficiently.
+Therefore, we focus on literature that explores strategies for utilizing existing bed resources more efficiently (Zamani et al.
+2024).
 Best et al.
@@
 As a follow-up, Izady and Mohamed (2021) consider a similar setting but introduce the option of “pooled wings,” which can accommodate patients from multiple specialties when dedicated wings reach capacity.
-Optimization techniques are primarily employed in this type of strategic planning to maximize inpatient bed utilization.
 Song et al.
@@
 However, our inpatient overflow MDP problem presents unique challenges that make the direct application of PPO algorithms difficult, specifically due to the large action space and the time-varying, periodic long-run average setting.
-In conventional PPO algorithms, the policy space is directly parameterized by an NN, with the output size corresponding to the number of possible actions – typically small in other applications (e.g., Schulman et al.
-(2017) and Dai and Gluzman (2022) focus on smaller action spaces).
+In conventional PPO algorithms, the policy space is directly parameterized by a Neural Network (NN), with the output size corresponding to the number of possible actions, typically a small number (e.g., Schulman et al.
+(2017), Dai and Gluzman (2022)).
 However, in our problem, the action space is significantly larger, making it computationally prohibitive to train the policy network using traditional parameterization methods.
-To our knowledge, there are no established examples in the literature where PPO has been successfully applied to such large action spaces, except for Feng et al.
-(2021), which showcase an effective use of action decomposition in ride-hailing applications.
-However, their setting is finite-horizon, whereas the framework we develop in this paper tackles the more complex infinite-horizon problem.
-Additionally, to account for the periodic nature of our infinite-horizon MDP, we introduce a novel partially-shared network structure that outputs actions at different epochs throughout the day.
-This design is conceptually related to specialized architectures in federated learning (Arivazhagan et al.
-2019, Kulkarni et al.
-2020), which address statistical heterogeneity, a different problem from ours.
+Feng et al.
+(2021) demonstrates the effectiveness of combining action decomposition with PPO for ride-hailing, but only in a finite-horizon setting.
+While PPO has been successfully applied to infinite-horizon MDPs (Dai and Gluzman 2022), to our knowledge, there are no established examples of its application to problems combining both large action spaces and infinite-horizon objectives.
+Our work addresses both challenges simultaneously, handling large action spaces in a time-varying, infinite-horizon environment.
+The study of atomic-action-based PPO is actively growing.
+While this paper was under review, concurrent work by Huo (2025) and Dai et al.
+(2025a) have demonstrated the effectiveness of this framework in data switch scheduling, and in battery charging and electric vehicle scheduling for robo-taxi systems, respectively; the latter introduced the term Atomic-PPO.
+Motivated by the empirical success, Dai et al.
+(2025b) further provides a theoretical justification for atomic action decomposition.
 3 Model Formulation
@@
 We formulate the overflow decision problem as an MDP in Section 3.2.
-Table 5 in Appendix 8 summarizes all the notations introduced in the model formulations and those used in later sections.
+Appendix 8 summarizes all the notations.
 3.1 Parallel-server System
@@
 This latter scenario is called an overflow assignment, and the non-primary pool that can serve class $j$ customers is referred to as an overflow pool for class $j$.
-[…]
+Arrival process.
+The underlying system evolves on a continuous-time scale.
+We assume that the arrival process for each customer class follows a periodic Poisson process, with the associated arrival rate function $\lambda_{j}(t)$ being periodic with a period of $T$ for each $j\in\mathcal{J}$.
+For ease of exposition, we set $T=1$ day to capture the time-varying pattern within a day.
+The corresponding daily arrival rate for class $j$ is given by $\Lambda_{j}=\int_{0}^{1}\lambda_{j}(t)dt$.
 Departure process.
@@
 We consider a discrete-time setup, where the decision maker (e.g., hospital manager) makes overflow decisions at $m$ predefined time epochs each day (e.g., hourly), denoted as $t_{0},t_{1},\dots,t_{m-1}$ for day $t$.
-In the rest of this section, we define the state, action, cost, policy, and objective of the MDP.
 State.
@@
   • $X_{j}(t_{k})$, $j=1,\dots,J$, denotes the customer count for pool $j$.
-This customer count includes both the number of customers waiting in buffer $j$ and those in service in pool $j$; we show below on how to recover each part.
-The former part (waiting count) only includes class $j$ customers, while the latter (in-service count) may include customers from other classes due to overflow assignments.
+This count includes both the number of customers waiting in buffer $j$ and those in service in pool $j$; we show below on how to recover each part.
+The former (waiting count) only includes class $j$ customers, while the latter (in-service count) may include customers from other classes due to overflow assignments.
   • $Y_{j}(t_{k})$, $j=1,\dots,J$, denotes the to-depart count in pool $j$.
-This quantity represents the number of customers expected to depart from pool $j$ between the present time and the end of the current day.
+This quantity represents the number of customers anticipated to depart from pool $j$ between the present time and the end of the current day.
 Tracking this quantity is essential for the two-time-scale inpatient departure scheme, where the “planned discharge” information is used to inform overflow decisions.
@@
 We use $q_{i}=(x_{i}-N_{i})\vee 0$ to denote this queue length for class $i$ (buffer $i$).
+Holding overflow decisions until predefined decision epochs allows managers to use up-to-date system information for better coordination.
+This practice is common in hospital assignment settings (Bertsimas and Pauphilet 2024, Han et al.
+2021) and has also been shown to improve decision quality in other operational contexts by allowing information to accumulate (Xie et al.
+2025).
 State transition.
@@
 After an action is taken, the resulting state is denoted as $s^{+}$, i.e., the post-action state.
-Given an action $f={f_{i,j}}$, the transition from $s$ to $s^{+}$ is as follows:
-$x^{+}_{j}$ $=x_{j}+\sum_{i=1,i\neq j}^{J}f_{i,j}-\sum_{\ell=1,\ell\neq j}^{J}f_{j,\ell},\quad j=1,\dots,J;$ (2)
-$y_{j}^{+}$ $=y_{j},\quad j=1,\dots,J;$
-$h^{+}$ $=h.$
-We denote the state, after the exogenous events (random arrivals and departures) occurring between the current epoch $t_{k}$ and the next epoch $t_{k+1}$, as $s^{\prime}$.
-The corresponding transition probability from state $s$ to state $s^{\prime}$, through $s^{+}$ under action $f$, is denoted as $p(s^{\prime}|s,f)$.
+We further denote by $s^{\prime}$ the state after the exogenous events (random arrivals and departures) occurring between the current epoch $t_{k}$ and the next epoch $t_{k+1}$.
+The transition from $s$ to $s^{+}$ is the same across all epochs (adding overflow-in patients and subtracting overflow-out patients), whereas the transition from $s^{+}$ to $s^{\prime}$ differs between the midnight epoch ($k=0$) and non-midnight epochs ($k\neq 0$).
+For non-midnight epochs, given an action $f={f_{i,j}}$, the transition from $s$ to $s^{\prime}$ follows:
+$x_{j}^{\prime}$ $=x_{j}^{+}+a_{j}-d_{j}=\left(x_{j}+\sum_{i=1,i\neq j}^{J}f_{i,j}-\sum_{k=1,k\neq j}^{J}f_{j,k}\right)+a_{j}-d_{j};\quad j=1,\dots,J,$ (2)
+$y_{j}^{\prime}$ $=y_{j}-d_{j};\quad j=1,\dots,J,$
+$h^{\prime}$ $=(h+1)\mod m,$
+where $a_{j}$ and $d_{j}$ denote, respectively, the number of class-$j$ arrivals and departures from pool $j$ between the two epochs.
+The distributions of arrivals and departures are specified in Appendix 9.
+For the midnight epoch, the main difference in the transition is in the to-depart count, with $y^{\prime}_{j}\sim\text{Bin}(\min\{x_{j},N_{j}\},\mu_{j})$ for $j=1,\dots,J$, meaning that each customer currently in service independently departs on the next day with probability $\mu_{j}$.
+We denote the transition probability from state $s$ to state $s^{\prime}$, through $s^{+}$ under action $f$, is denoted as $p(s^{\prime}|s,f)$.
 The detailed calculation of these transition probabilities is provided in Appendix 9.
@@
 $g(s,f)=\sum_{j=1}^{J}C_{j}q_{j}^{+}+\sum_{i=1}^{J}\sum_{j=1,j\neq i}^{J}B_{i,j}f_{i,j},$
-where $q_{j}^{+}=(x_{j}^{+}-N_{j})\vee 0$ denotes the post-action queue length for class $j$, and the transition to the post-action state follows (2).
-This cost setup is designed to identify a policy that strikes a balance between the negative effects of waiting and the undesirability of overflow assignments.
+where $q_{j}^{+}=(x_{j}^{+}-N_{j})\vee 0$ denotes the post-action queue length for class $j$.
 Policy and objective.
@@
 We focus on all feasible stationary Markovian policies, and we denote the policy space as $\Pi$.
-For a policy $\pi\in\Pi$, we use $\pi(f|s)$ to denote the probability of selecting a feasible action $f$ in a given state $s$ (where deterministic policies are included via using the Dirac measure).
+For a policy $\pi\in\Pi$, we use $\pi(f\mid s)$ to denote the probability of selecting a feasible action $f$ in a given state $s$ (where deterministic policies are included via using the Dirac measure).
 The resulting stochastic process $\{S(t_{k})\}$ forms a discrete-time Markov Chain (DTMC) with a one-epoch transition probability from state $s$ to $s^{\prime}$ as $p_{\pi}(s^{\prime}|s)=\mathbb{E}_{f\sim\pi(\cdot|s)}p(s^{\prime}|s,f)$.
@@
 In addition, although we assume a linear cost function in this paper, our framework can handle more complex cost structures, such as polynomial costs or costs related to service levels (e.g., the proportion of patients waiting more than 6 hours).
-See the appendix of Dai and Shi (2019) for an example of modifying the cost to account for service-level metrics.
+See the appendix of Dai and Shi (2019) for an example to account for service-level metrics.
 We note that these adjustments do not alter the core elements of our problem, and our algorithmic framework is both flexible and general enough to accommodate them.
-For ease of exposition, throughout the remainder of the paper, we will focus on the current baseline setting, where primary assignments are handled automatically, departures are pool-dependent, no transfer back of overflow customers, and the cost is linear.
+For ease of exposition, we will focus on the current baseline setting, where primary assignments are handled automatically, departures are pool-dependent, no transfer back of overflow customers, and the cost is linear.
 4 Atomic Actions and Proximal Policy Optimization
 One of the core challenges in solving the MDP model developed in Section 3.2 is the combinatorial explosion of the action space.
-When assigning servers for multiple customers simultaneously, the number of possible combinations grows exponentially with the number of customer classes and server pools.
 Specifically, when there are enough idle servers in the system, the action space has a size of $\prod_{j=1}^{J}\left(\begin{array}[]{c}q_{j}+w_{j}\\ w_{j}\end{array}\right)$, where $q_{j}$ is the number of waiting customers in class $j$ and $w_{j}$ is the number of overflow pools for class $j$ that have idle servers.
 To illustrate, in a twenty-pool system with each server-pool having a capacity of 60 and a state $(x_{1},x_{2},x_{3},\dots,x_{20})=(65,63,62,50,50,50,\dots,50)$, the number of potential ways to assign 10 waiting customers can exceed $10^{9}$.
-In other words, even if the value function is computed (either exactly or via approximation), brute-force searching through all possible actions to find the one with the lowest cost-to-go is infeasible.
-To overcome this curse of dimensionality, we use atomic actions in combination with PPO to directly optimize the policy, which forms the key foundation for our algorithm’s scalability.
+Even if the value function is computed (either exactly or via approximation), brute-force searching through all possible actions to find the one with the lowest cost-to-go is infeasible.
+To overcome this curse of dimensionality, we use atomic actions in combination with PPO that form the key foundation for our algorithm’s scalability.
 In Section 4.1, we introduce the atomic action and policy setup.
-In Section 4.2, we present the framework of the PPO algorithm.
+In Section 4.2, we present the PPO framework.
 4.1 Action Space Reduction: Atomic Actions
@@
 The action space for each $a^{n}$ is simply $\mathcal{J}=\{1,\dots,J\}$.
-If the class of the $n$th customer is $j$, then $a^{n}=j$ indicates the customer will continue waiting until the next decision epoch; otherwise, the customer is assigned to an overflow pool (see more details on feasible overflow assignment in Section 4.1.2).
-$s^{1}=s$$s^{\prime}$$\textbf{a}=(a^{1},\dots,a^{q})$$s^{2}$$s^{3}$$s^{q+1}$$\sigma^{1}(s),a^{1}$$\sigma^{2}(s),a^{2}$rrivalDepartureTime $t_{k}$ime $t_{k+1}$
+If the class of the $n$th customer is $j$, then $a^{n}=j$ indicates the customer will continue waiting until the next decision epoch; otherwise, the customer is assigned to an overflow pool (see details on feasible overflow assignment in Section 4.1.2).
+$^{1}=sss^{\prime}$$\textbf{a}=(a^{1},\dots,a^{q})$$s^{2}$$s^{3}$$s^{q+1}$$\sigma^{1}(s),a^{1}$$\sigma^{2}(s),a^{2}$ArrivalDepartureTime $t_{k}$Time $t_{k+1}$
 Figure 1: Illustration of the atomic decision process.
@@
 Starting from the first customer, who is from class $\sigma^{1}(s)$, we execute the atomic action $a^{1}$ for this customer, then we update the state according to $a^{1}$ and sequentially move to the second customer, so on and forth till we execute the atomic action $a^{q}$ for the last waiting customer, who is from class $\sigma^{q}(s)$.
-The atomic action $a^{n}$ for the $n$th customer is chosen based on the atomic state, $s^{n}$, which is observed after assigning the $(n-1)$th customer.
+The atomic action $a^{n}$ for the $n$th customer is chosen based on the atomic state, $s^{n}$, which is observed after assigning the (n1)(n-1)th customer.
 This state has the same structure as the one introduced in Section 3.2, i.e., $s^{n}=(x^{n}_{1},\dots,x^{n}_{J},y_{1},\dots,y_{J},h)$.
@@
 This reduction forms the most critical foundation for developing an algorithm that scales effectively with the number of customer classes and server pools.
-For clarity, and to distinguish between the MDP state and action defined in Section 3.2 and those used for the atomic decision process here, we will refer to the state and action in Section 3.2 as the system-level state and action throughout the rest of the paper.
-Meanwhile, we will refer to the state and action in the atomic decision process as the atomic state and action.
+For clarity, and to distinguish between the MDP state and action defined in Section 3.2 and the ones used for the atomic decision process here, we will refer to those in Section 3.2 as the system-level state and action throughout the rest of the paper.
+Meanwhile, we will refer to the ones in the atomic decision process as the atomic state and action.
 It is important to note that the system-level and atomic states and actions are closely related.
 Specifically, the initial atomic state $s^{1}$ corresponds to the system-level (pre-action) state $s$, and after all atomic actions are executed, the final atomic state $s^{q+1}$ corresponds to the system-level post-action state $s^{+}$.
-Additionally, the sum of the atomic actions for the inflow and outflow of class $j$ customers can be used to recover the system-level actions.
-That is, for any given pair $(\boldsymbol{\sigma}(s),\textbf{a})$, we can recover the system-level action $f_{i,j}$ via $\sum_{n=1}^{q}\mathbbm{1}(a^{n}=j,\sigma^{n}(s)=i)$.
+For any given pair $(\boldsymbol{\sigma}(s),\textbf{a})$, we can recover the system-level action $f_{i,j}$ via $\sum_{n=1}^{q}\mathbbm{1}(a^{n}=j,\sigma^{n}(s)=i)$.
+Theoretical support for the proposed atomic action approach comes from a contemporary paper Dai et al.
+(2025b).
+Their Theorems 1 and 2 show that, in certain stochastic processing networks, (i) decomposing joint actions into atomic steps incurs no loss in optimality, and (ii) even enforcing step-independent policies (each atomic action follows the same rule) still achieves the optimal long-run average reward.
+This step-independent structure mirrors our setup in Section 4.1, where each atomic action is sampled from the same policy $\pi(\cdot\mid s^{\ell})$ regardless of step index $\ell$.
 4.1.1 Randomized Atomic Action
 In this paper, we use randomized policies to prescribe the atomic actions.
-We will first outline the specifics of this randomization and then discuss its advantages over conventional deterministic policies in Section 4.1.2.
+We will first outline the specifics of this randomization and then discuss its advantages over conventional deterministic policies at the end of Section 4.2.
 For a given atomic state $s^{n}$, we define a probability distribution over all possible atomic actions for each customer class.
 We denote this distribution for a class $i$ customer, given $s^{n}$, as $\kappa(\cdot|s^{n},i)$, and the set of distributions as $\kappa=\{\kappa(\cdot|s^{n},i),i\in\mathcal{J}\}$.
-Consequently, starting from the initial state $s^{1}=s$ (the system-level pre-action state) and following a given order $\boldsymbol{\sigma}(s)$, the probability of selecting a sequence of atomic actions $\textbf{a}=\{a^{1},\dots,a^{q}\}$ is given by $\prod_{n=1}^{q}\kappa(a^{n}|s^{n},\sigma^{n}(s))$, where $(s^{1},\dots,s^{q})$ denotes the corresponding atomic state sequence generated within the atomic decision process, as specified in Equation (4).
+Consequently, starting from the initial state $s^{1}=s$ (the system-level pre-action state) and following a given order $\boldsymbol{\sigma}(s)=\{\sigma^{1}(s),\dots,\sigma^{q}(s)\}$ for $q$ waiting customers, the probability of selecting a sequence of atomic actions $\textbf{a}=\{a^{1},\dots,a^{q}\}$ is given by $\prod_{n=1}^{q}\kappa(a^{n}|s^{n},\sigma^{n}(s))$, where $(s^{1},\dots,s^{q})$ denotes the corresponding atomic state sequence generated within the atomic decision process, as specified in Equation (4).
 According to the correspondence between atomic and system-level actions, the probability of taking a system-level action $f$ under the given distributions $\kappa$ can be expressed as:
-$\pi(f|s)=$ $\sum_{\textbf{a}}\prod_{n=1}^{q}\kappa(a^{n}|s^{n},\sigma^{n}(s))\cdot\mathbbm{1}\Big{(}\sum_{n=1}^{q}\mathbbm{1}(a^{n}=j,\sigma^{n}(s)=i)=f_{i,j},\forall i,j\Big{)}.$ (5)
-To simplify notation and align with the broader PPO algorithm literature, we use the system-level policy $\pi=\{\pi(f|s)\}$ in the rest of the paper when we refer to a policy.
+$\pi(f\mid s)=$ $\sum_{\textbf{a}}\prod_{n=1}^{q}\kappa(a^{n}|s^{n},\sigma^{n}(s))\cdot\mathbbm{1}\Big(\sum_{n=1}^{q}\mathbbm{1}(a^{n}=j,\sigma^{n}(s)=i)=f_{i,j},\forall i,j\Big).$ (5)
+To simplify notation and align with the broader PPO algorithm literature, we use the system-level policy $\pi=\{\pi(f\mid s)\}$ in the rest of the paper when we refer to a policy.
 However, it is important to clarify that this system-level policy is driven by the underlying atomic-level probability distributions.
@@
 Our goal is to optimize $\kappa$ in such a way that the induced system-level policy $\pi$ achieves the performance objective in (3).
-Moreover, in Section 5.3, we will introduce a “batching” setup that simplifies the calculation of (5) and eliminates the dependence on the specific processing order $\boldsymbol{\sigma}$.
-This will establish a more direct connection between the system-level policy $\pi$ and the atomic-level distributions $\kappa$.
 We discuss how the randomized policy can be practically applied in decision-making in Section 6.5.
@@
 $\kappa_{\theta}\big(a^{n}=j|s^{n},\sigma^{n}(s)=i\big)=\frac{\text{exp}(g_{i,j}(s^{n}|\theta))}{\sum_{k=1}^{J}\text{exp}(g_{i,k}(s^{n}|\theta))},$ (6)
-where $g_{i,j}(s|\theta)$ is one of the $J\times J$ output neurons of a dense NN with input $s$ and parameters $\theta$. In the simplest setting, we can use a NN with no hidden layers, i.e.,
-$g_{i,j}\big{(}s|\theta=(w_{i,j},b_{i,j},i,j\in\mathcal{J})\big{)}=w_{i,j}\cdot s+b_{i,j},$
-which is equivalent to a multi-class logistic regression model. For further details on NN parameterization, see Section 5.1.
-Note that Equation (6) could assign a positive probability to non-feasible atomic actions, such as when the selected server pool is full or cannot serve class $j$.
+where $g_{i,j}(s|\theta)$ is one of the $J\times J$ output neurons of a dense NN with input $s$ and parameters $\theta$.
+Equation (6) could assign a positive probability to non-feasible atomic actions, such as when the selected server pool is full or cannot serve class $j$.
 To ensure the feasibility of the prescribed atomic action, we set $\kappa_{\theta}(a^{n}|s^{n},\sigma^{n}(s))=0$ for $a^{n}\not\in\mathcal{A}(s^{n},\sigma^{n}(s))$ (i.e., force the probability of non-feasible actions to 0), and then rescale the remaining probabilities to sum to 1.
@@
 4.2 Finding Optimal Policy: Proximal Policy Optimization
-The randomized policy framework enables the use of policy gradient methods to learn the parameter $\theta$.
-In this paper, we employ the state-of-the-art policy gradient method, PPO, with custom modifications tailored to our specific queueing network context.
+We employ the state-of-the-art policy gradient method, PPO, to learn the parameter $\theta$ with custom modifications tailored to our specific queueing network context.
 We provide an overview of the PPO framework in this section; we defer the details of our tailored modifications to Section 5.
@@
 (2017) suggests minimizing the following PPO objective:
-$L(\theta):=\mathop{\mathbb{E}}\limits_{\begin{subarray}{c}s\sim\mu_{\eta}\\ f\sim\pi_{\eta}(\cdot|s)\end{subarray}}\max\left\{r_{\theta,\eta}(f|s)A_{\eta}(s,f),\quad\text{clip}\left(r_{\theta,\eta}(f|s),1-\epsilon,1+\epsilon\right)A_{\eta}(s,f)\right\}.$ (9)
+$L(\theta):=\mathop{\mathbb{E}}\limits_{\begin{subarray}{c}s\sim\mu_{\eta}\\ f\sim\pi_{\eta}(\cdot|s)\end{subarray}}\max\left\{r_{\theta,\eta}(f\mid s)A_{\eta}(s,f),\quad\text{clip}\left(r_{\theta,\eta}(f\mid s),1-\epsilon,1+\epsilon\right)A_{\eta}(s,f)\right\}.$ (9)
 Here, the distribution $\mu_{\eta}$ denotes the stationary distribution under policy $\pi_{\eta}$; $r_{\theta,\eta}$ denotes the probability ratio between $\pi_{\theta}$ and $\pi_{\eta}$:
-$r_{\theta,\eta}(f|s):=\frac{\pi_{\theta}(f|s)}{\pi_{\eta}(f|s)},$ (10)
-where $\pi_{\theta}(f|s)$ and $\pi_{\eta}(f|s)$ are defined as in (5), with $\kappa$ replaced by $\kappa_{\theta}$ and $\kappa_{\eta}$, respectively; and the clipped function
+$r_{\theta,\eta}(f\mid s):=\frac{\pi_{\theta}(f\mid s)}{\pi_{\eta}(f\mid s)},$ (10)
+where $\pi_{\theta}(f\mid s)$ and $\pi_{\eta}(f\mid s)$ are defined as in (5), with $\kappa$ replaced by $\kappa_{\theta}$ and $\kappa_{\eta}$, respectively; and the clipped function
 $\text{clip}(x,1-\epsilon,1+\epsilon):=\left\{\begin{array}[]{cc}1-\epsilon,&\text{if }x<1-\epsilon\\ 1+\epsilon,&\text{if }x>1+\epsilon\\ x,&\text{otherwise}\end{array}\right.$.
-The acceptable difference $\epsilon$ is a hyper-parameter that can be tuned for better algorithm performance.
+The difference $\epsilon$ is a hyper-parameter that can be tuned for better algorithm performance.
 For the theoretical guarantee, we establish Proposition 1, which extends the performance improvement guarantee of PPO from the time-stationary, long-run average cost setting in Dai and Gluzman (2022) to our more complex periodic environment.
-See the complete statement of Proposition 1 (alongside needed technical conditions) in Appendix 10 and its proof in Appendix 13.
+See the complete statement of Proposition 1 (alongside needed technical conditions) in Appendix 10.
 The key challenge in the proof lies in the periodicity of state transitions, which makes directly applying previous results inapplicable.
 To address this, we introduce a new framework using daily Markov chains (MCs), where one period (e.g., one day) is treated as a single transition step.
-This transformation allows us to redefine the objective function in a way that ensures aperiodicity and enables the application of previous theoretical framework.
+This transformation allows us to redefine the objective function in a way that ensures aperiodicity to use previous theoretical frameworks.
 Proposition 1 establishes that the change in long-run average cost can be decomposed into two terms: one that dominates the improvement and another that diminishes at a faster rate.
@@
 In implementation, we estimate the objective function (9) using simulation data samples, i.e.,
-$\hat{L}(\theta,\mathcal{D}_{\eta}^{T},\hat{\textbf{A}}_{\eta}(\mathcal{D}_{\eta}^{T}))=\frac{1}{Tm}\sum_{t=1}^{T}\sum_{k=0}^{m-1}\max$$\left\{r_{\theta,\eta}(f(t_{k})|s(t_{k}))\hat{A}_{\eta}(s(t_{k}),f(t_{k})),\right.$ (11)
-$\left.\text{clip}\left(r_{\theta,\eta}(f(t_{k})|s(t_{k})),1-\epsilon,1+\epsilon\right)\hat{A}_{\eta}(s(t_{k}),f(t_{k}))\right\},$
-where the dataset $\mathcal{D}_{\eta}^{T}=\left\{\big{(}s(t_{k}),f(t_{k}),g(t_{k})\big{)}_{k=0}^{m-1}\right\}_{t=1}^{T}$ contains $T$ days of simulation data generated under policy $\pi_{\eta}$, and the vector $\hat{\textbf{A}}_{\eta}(\mathcal{D}_{\eta}^{T})=\left\{\left(\hat{A}_{\eta}(s(t_{k}),f(t_{k}))\right)_{k=0}^{m-1}\right\}_{t=1}^{T}$ consists of the estimated advantage functions for each decision epoch within the $T$ days using the collected data $\mathcal{D}_{\eta}^{T}$.
-Note that the constant factor $\frac{1}{Tm}$ in $\hat{L}$ can be disregarded without affecting the objective optimization.
+$\hat{L}(\theta,\mathcal{D}_{\eta}^{T},\hat{\textbf{A}}_{\eta}(\mathcal{D}_{\eta}^{T}))=\frac{1}{Tm}\sum_{t=1}^{T}\sum_{k=0}^{m-1}\max$ $\left\{\hat{r}_{\theta,\eta}(f(t_{k})|s(t_{k}))\hat{A}_{\eta}(s(t_{k}),f(t_{k})),\right$.
+(11)
+$\left.\text{clip}\left(\hat{r}_{\theta,\eta}(f(t_{k})|s(t_{k})),1-\epsilon,1+\epsilon\right)\hat{A}_{\eta}(s(t_{k}),f(t_{k}))\right\},$
+where the dataset $\mathcal{D}_{\eta}^{T}$ contains $T$ days of simulation data generated under policy $\pi_{\eta}$ (including the atomic states, atomic actions and order of patients within the simulation of sequential making process for each day), and the vector $\hat{\textbf{A}}_{\eta}(\mathcal{D}_{\eta}^{T})=\left\{\left(\hat{A}_{\eta}(s(t_{k}),f(t_{k}))\right)_{k=0}^{m-1}\right\}_{t=1}^{T}$ consists of the estimated advantage functions for each decision epoch within the $T$ days using the collected data $\mathcal{D}_{\eta}^{T}$.
 The detailed steps of the PPO algorithm are given in Algorithm 1.
@@
 Input: Policy $\pi_{\eta}$
-1:  for policy iteration $r=0,1,2,...,R-1$  do
-2:     Run policy $\pi_{\eta}$ for $T$ days and collect simulation dataset $\mathcal{D}_{\eta}^{T}=\left\{\big{(}s(t_{k}),f(t_{k}),g(t_{k})\big{)}_{k=0}^{m-1}\right\}_{t=1}^{T}$.
+1:for policy iteration r=0,1,2,,R1r=0,1,2,...,R-1 do
+2:  Run policy $\pi_{\eta}$ for $T$ days and collect simulation dataset $\mathcal{D}_{\eta}^{T}$.
 3:  Estimate the advantage functions $\hat{\textbf{A}}_{\eta}(\mathcal{D}_{\eta}^{T})$.
@@
 6:end for
+Computing $\pi(f\mid s)$ in (5) is computational expensive due to the required summation.
+In implementation, we approximate it using the product $\prod_{n=1}^{q(t_{k})}\kappa\big(a^{n}(t_{k})\mid s^{n}(t_{k}),\sigma^{n}(t_{k})\big)$, evaluated along the realized processing order in simulation; correspondingly, we compute an approximate policy ratio $\hat{r}_{\theta,\eta}$.
+In Section 5.3, we further introduce a batching scheme that eliminates order dependence and yields a simpler, more direct mapping between the system-level policy $\pi$ and the atomic-level distributions $\kappa$.
 We conclude this section with two important remarks.
-First, our policy parameterization differs from the conventional ones used in PPO framework, which would directly parameterize a randomized policy over the entire combinatorial system-level action space.
-This action space has a size of $\prod_{j=1}^{J}\left(\begin{array}[]{c}q_{j}+w_{j}\\ w_{j}\end{array}\right)$, which means that a conventional policy network would require an extremely large output layer, making training highly challenging.
-The key element for our algorithm is the combination of atomic actions and the randomized policy.
-Second, performance improvement using the PPO algorithm is ensured when policy updates are conservative (see Appendix 10).
-The atomic action probabilities are parameterized using continuous functions that output distributions over small action spaces, allowing incremental updates that are scalable given the relatively small parameter size.
-This would not be feasible if we use the deterministic policy setup.
-We have also conducted numerical experiments that demonstrate the superior performance of our randomized approach compared to deterministic ones.
-The numerical results highlight two major advantages of randomized atomic actions: (i) avoiding chattering – randomized policies integrate well with PPO, which prevent policy “chattering” in discrete-action spaces; (ii) randomized actions naturally allow for exploration, which improve the speed of convergence to a “good” final policy.
-We discuss how to use the randomized policy in practice in Section 6.5.
-5 PPO Design: Tailoring for Periodicity in MDP
+First, our atomic policy parameterization differs from the conventional ones used in PPO framework, which directly parameterize a randomized policy over the entire system-level action space and would otherwise require an intractably large policy network.
+Second, performance improvement in PPO relies on conservative policy updates (Proposition 1 in Appendix 10), which is infeasible under the deterministic policy setup.
+In particular, while atomic actions alone reduce the action space, combining them with traditional value-based ADP leads to severe policy “chattering.” This phenomenon arises because, when value functions are approximated rather than evaluated exactly, greedy policy updates may overreact to estimation errors, causing oscillations between improving and worsening policies (Bertsekas and others 2011).
+PPO’s proximal updates mitigate this issue.
+In other words, the combination of atomic actions and PPO is critical; see empirical results in Section 6.
+5 Tailoring Design for Periodicity in MDP
 A salient feature of our problem setting is the intrinsic time-dependency.
@@
 The partially-shared network excludes the epoch index from input but uses different blocks of output neurons for actions of different epochs.
-The parameters of the first several layers are shared across epochs and while those of the later layers are specific to each epoch.
   • The fully-connected network structure includes the epoch index as part of its input, and generates atomic action probabilities of different epochs with the same set of NN parameters.
@@
     where $g_{i,j}$ is one of the $J\times J$ output neurons of the fully-connected network.
-This design is the most straightforward way to handle time-dependency – the network uses the same set of parameters to generate outputs for different epochs, which captures temporal differences by incorporating time as part of the input.
+This design uses the same set of parameters to generate outputs for different epochs, which captures temporal differences by incorporating time as part of the input.
   • The fully-separate network structure uses $m$ independent networks for each of the $m$ different epochs, so the epoch index does not need to be part of the input.
@@
 The fully-connected and fully-separate designs each have their pros and cons, while the partially-shared policy network combines their strengths.
-Specifically, the fully-connected network design uses the same set of parameters across all time epochs, naturally capturing similarities between epochs and enabling training with data from all time epochs.
+The fully-connected network design uses the same set of parameters across all time epochs, naturally capturing similarities between epochs and enabling training with data from all time epochs.
 However, since the policy differences between epochs can be highly complex, a more sophisticated network (e.g., with additional hidden layers) may be needed to capture these nuances and hence more training data.
@@
 The partially-shared network design strikes a balance: shared parameters in the early layers capture commonalities between epochs for efficient training, while the separated last layer adapts to epoch-specific policy outputs.
-In the case study of a ten-pool system, the partially-shared design requires only a few thousand simulation data samples for training, yet outperforms the other two designs that need 100 times more data.
-We will provide a detailed comparison of the sample efficiency of these three designs in Section 6.4.
+In the case study of a ten-pool system, the partially-shared design requires only a few thousand simulation data samples for training, yet outperforms the other two designs that need 100 times more data; see more in Section 6.4.
 5.2 Policy Evaluation: Linear Value Function Approximation
 In the standard PPO implementation, a separate neural network, known as the value network, is used alongside the policy network to approximate value functions.
-The value network is trained with targets estimated via Monte Carlo evaluation.
-These estimates can have high variance in long-run average settings, which require large amounts of data to achieve accurate policy evaluation.
-To address this, we adopt a linear value function approximation and leverage the queueing system structure to design proper basis functions.
-This approach mitigates variance-related inefficiencies and contrasts with the black-box nature of NN-based value estimation.
-Additionally, our approach extends to other operations problems with similar structural properties.
+The value network is trained with targets estimated via Monte Carlo evaluation, which can have high variance in long-run average settings.
+In this paper, we adopt a linear value function approximation and leverage the queueing system structure to design proper basis functions.
+To train the value approximation, we use the Least-Squares Temporal Difference (LSTD) method proposed by Bradtke and Barto (1996), which provides more stable estimates in tabular and linear function approximation settings, as shown in Boyan (1999) and Lagoudakis and Parr (2003).
 The key to accurate linear value function approximations lies in the choice of basis functions. Our proposed basis functions are:
@@
 Each $\hat{v}_{\pi,j}(s_{j})$ relies only on pool-$j$’s state information, $s_{j}=(x_{j},y_{j},h)$, and estimates the relative value function based on the holding and overflow cost components for class-$j$ customers.
-Specifically, under the randomized atomic policy $\pi$, we decompose the original $J$-pool system into $J$ single-pool systems.
-For each decomposed pool-$j$ system, the relative value function $\hat{v}_{\pi,j}$ is solved via the corresponding post-action Poisson equation
+Specifically, under the randomized atomic policy $\pi$, we decompose the original $J$-pool system into $J$ single-pool systems, and the relative value function $\hat{v}_{\pi,j}$ for the decomposed pool-$j$ system is solved via the post-action Poisson equation
 $\hat{v}_{\pi,j}(s_{j})=C_{j}q_{j}-\hat{\gamma}_{\pi,j}+\mathbb{E}\left[\sum_{k=1}^{J}B_{j,k}\bar{f}_{j,k}+\hat{v}_{\pi,j}(s_{j}^{\prime})\right]$ (13)
-for a given post-action state $s_{j}$ (we drop the superscript $+$ here for simplicity).
+for a given post-action state $s_{j}$ (the superscript ++ dropped for simplicity).
 Here, $C_{j}q_{j}$ is the post-action holding cost with $q_{j}=\max(s_{j}-N_{j},0)$, $\hat{\gamma}_{\pi,j}$ is the average cost of the decomposed pool-$j$ system, and the cost-to-go term consists of: (a) the expected overflow cost $\mathbb{E}[\sum_{k=1}^{J}B_{j,k}\bar{f}_{j,k}]$, where $\bar{f}_{j,k}$ approximates the overflow assignments from class $j$ to other pools, and (b) the expected value function at the next post-action state $\mathbb{E}\left[\hat{v}_{\pi,j}(s’_{j})\right]$.
@@
 Given the arrivals are Poisson and leveraging the Poisson thinning property, the outflows $\bar{f}_{j,k}$ and inflows $\bar{f}_{i,j}$ can be approximated as Poisson distributions with rates $\lambda_{j}\bar{\kappa}_{j}(k|s_{j},j)$ and $\lambda_{i}\bar{\kappa}_{i}(j|s_{j},i)$, respectively.
-This formulation allows efficient evaluation of both (a) and (b).
-Details on computing $\bar{\kappa}_{j}$, deriving $\hat{v}_{\pi,j}$, and formulating (13) using post-action states are provided in Appendix 11.
-The feasibility of the pool-wise decomposition relies on the randomized atomic policy.
-The continuous nature of overflow probabilities under this policy allows us to generate good approximations of the $f_{i,j}$ values.
-In contrast, a deterministic policy results in discrete actions with more drastic variations across states, making it harder to approximate overflow proportions effectively.
-This distinction makes our new basis function, $V_{d}$, specifically suited for the PPO framework and different from the $V_{s}$ basis function proposed in Dai and Shi (2019), which was designed for deterministic, discrete-action spaces.
-In Section 6.4, we compare the performance of $V_{d}$ and $V_{s}$ alongside other polynomial bases.
-Our results demonstrate that $V_{d}$ improves sample efficiency and value function approximation accuracy, which ultimately leads to better policy performance.
+Details on computing $\bar{\kappa}_{j}$ and deriving $\hat{v}_{\pi,j}$ are provided in Appendix 11.
 5.3 Batching Setup and Other Implementation Details
 In the general atomic decision process, each of the atomic actions sequentially depends on the atomic states as illustrated in Figure 1.
-We propose a batching setup to simplify this process, which uses the same initial state $s=s^{1}$ for all atomic actions in a batch.
+We find that decision ordering has limited impact on overall performance; see Appendix 12.2.4.
+Motivated by this observation, we propose a batching setup to simplify the decision process, which uses the same initial state $s=s^{1}$ for all atomic actions in a batch.
 Specifically, given $s$ and a chosen customer class order $\boldsymbol{\sigma}(s)$, the probability of conducting atomic action sequence a becomes $\prod_{n=1}^{q}\kappa(a^{n}|s,\sigma^{n}(s))$.
@@
 Under the batching setup, for a given system-level state $s$ and any pre-determined rule for choosing customer order $\boldsymbol{\sigma}(s)$, the probability of taking action $f$ follows a multinomial distribution that is independent of the order $\boldsymbol{\sigma}(s)$, i.e.,
-$\pi(f|s)=\prod_{i=1}^{J}\frac{q_{i}!}{\prod_{j=1}^{J}f_{i,j}!}\prod_{j=1}^{J}\kappa(j|s,i)^{f_{i,j}}.$ (14)
+$\pi(f\mid s)=\prod_{i=1}^{J}\frac{q_{i}!}{\prod_{j=1}^{J}f_{i,j}!}\prod_{j=1}^{J}\kappa(j|s,i)^{f_{i,j}}.$ (14)
 Proof 5.2
 Proof. From (5), we have that
-$\pi(f|s)=$ $\sum_{\textbf{a}}\prod_{n=1}^{q}\kappa(a^{n}|s,\sigma^{n}(s))\cdot\mathbbm{1}\left(f_{i,j}=\sum_{n=1}^{q}\mathbbm{1}(a^{n}=j,\sigma^{n}(s)=i),i,j\in\mathcal{J}\right)$
+$\pi(f\mid s)=$ $\sum_{\textbf{a}}\prod_{n=1}^{q}\kappa(a^{n}|s,\sigma^{n}(s))\cdot\mathbbm{1}\left(f_{i,j}=\sum_{n=1}^{q}\mathbbm{1}(a^{n}=j,\sigma^{n}(s)=i),i,j\in\mathcal{J}\right)$
 $=$ $\left(\prod_{i=1}^{J}\prod_{j=1}^{J}\kappa(j|s,i)^{f_{i,j}}\right)\cdot\sum_{\textbf{a}}\mathbbm{1}\left(f_{i,j}=\sum_{n=1}^{q}\mathbbm{1}(a^{n}=j,\sigma^{n}(s)=i),i,j\in\mathcal{J}\right)$
@@
 $\square$
-This batching setup is new in our application context compared to other atomic action approaches (e.g., Feng et al.
-(2021)), as our system-level overflow decisions are made on a discrete-time basis with multiple customers to be processed in a batch.
 The first advantage of this batching approach is that the atomic policy for the same class of customers remains consistent within the same decision epoch, which ensures a degree of fairness regardless of the pre-selected processing order $\boldsymbol{\sigma}(s)$.
@@
 Unlike the original sequential atomic decision setup, we do not need to update each atomic state or compute the state-dependent atomic action probabilities for each individual customer.
-The third advantage lies in the great simplification of calculating the objective function (11) for PPO.
+The third advantage lies in the great simplification of evaluating (11).
 Specifically, recall that we need to compute the probability ratio
-$r_{\theta,\eta}(f|s)=\frac{\pi_{\theta}(f|s)}{\pi_{\eta}(f|s)}$
-at each decision epoch, where $\pi_{\theta}(f|s)$ and $\pi_{\eta}(f|s)$ are defined as in (5), with $\kappa$ replaced by $\kappa_{\theta}$ and $\kappa_{\eta}$, respectively.
-Under the original setup, getting this ratio is computationally intensive because it involves summing over all possible atomic action sequences $\boldsymbol{a}$ corresponding to the system-level action $f$ for a given customer order $\boldsymbol{\sigma}(s)$ (i.e., the summation over $\boldsymbol{a}$ part in (5)).
+$r_{\theta,\eta}(f\mid s)=\frac{\pi_{\theta}(f\mid s)}{\pi_{\eta}(f\mid s)}$
+at each decision epoch, where $\pi_{\theta}(f\mid s)$ and $\pi_{\eta}(f\mid s)$ are defined as in (5), with $\kappa$ replaced by $\kappa_{\theta}$ and $\kappa_{\eta}$, respectively.
+Directly calculating this ratio is computationally intensive because it involves summing over all possible atomic action sequences $\boldsymbol{a}$ corresponding to the system-level action $f$ for a given customer order $\boldsymbol{\sigma}(s)$ (i.e., the summation over $\boldsymbol{a}$ part in (5)).
 In contrast, under the batching setup, using (14) from Lemma 14, this ratio is greatly simplified to:
-$r_{\theta,\eta}(f|s)=\prod_{i=1}^{J}\prod_{j=1}^{J}\left(\frac{\kappa_{\theta}(j|s,i)}{\kappa_{\eta}(j|s,i)}\right)^{f_{i,j}},$
+$r_{\theta,\eta}(f\mid s)=\prod_{i=1}^{J}\prod_{j=1}^{J}\left(\frac{\kappa_{\theta}(j|s,i)}{\kappa_{\eta}(j|s,i)}\right)^{f_{i,j}},$ (15)
 since the combinatorial factor $\frac{q_{i}!}{\prod_{j=1}^{J}f_{i,j}!}$ is a constant for given $f_{i,j}$’s and cancels out in the ratio calculation.
 This also establishes a direct equivalence between the ratio of atomic action probabilities ($\kappa_{\theta}$, $\kappa_{\eta}$) and the ratio of the induced system-level policies ($\pi_{\theta}$, $\pi_{\eta}$).
-As in the original setup, it is possible that during implementation a sampled atomic action under the batching setup could become infeasible if the selected pool no longer has an available server or cannot accommodate the customer class.
+As in the original setup, it is also possible that a sampled action under the batching setup could become infeasible if the selected pool no longer has an available server or cannot accommodate the customer class.
 To handle this, we apply the same resampling technique to ensure a feasible action is selected.
@@
 The PPO algorithm iterates through three major modules: simulation data generation, policy evaluation, and policy improvement.
-For data generation, multiple actors (CPUs) run parallel simulations of equal length, and then the simulated data is concatenated into one long episode.
+For data generation, multiple actors (CPUs) run parallel simulations of equal length, and then the simulated data is concatenated into one long trajectory, which constitutes the entire training dataset.
 For policy evaluation, we compute the linear value function approximations and the estimated advantage functions.
@@
 In Section 6.2, we show that our PPO algorithm achieves similar performance to the ADP method from Dai and Shi (2019) – considered the “state-of-the-art” – while significantly reducing computational time in the five-pool baseline, the largest system handled by Dai and Shi (2019).
-In Section 6.3, we focus the analysis on ten- and twenty-pool systems, where our method significantly outperforms several benchmark policies.
-In Section 6.4, we examine the algorithm’s performance across different hyper-parameter settings and highlight the critical role of our atomic policy design and tailored components.
-Finally, in Section 6.5 we summarize key takeaways and offers practical recommendations for managing inpatient overflow.
+In Section 6.3, we show the scalability in ten- and twenty-pool systems.
+In Section 6.4, we examine the algorithm’s performance across different hyper-parameter settings.
+We also conduct a detailed ablation study to highlight the critical role of our atomic policy design and the value of each tailored component.
+Finally, in Section 6.5 we showcase that the policy from the algorithm output is interpretable, and we summarize key takeaways for managing inpatient overflow.
 6.1 Model Settings
@@
 We set the baseline values at $B=(30,35)$ for overflow costs and $C=6$ for holding costs.
-These cost parameters act as “tuning knobs,” which allow managers to adjust them to generate Pareto curves and balance congestion with overflow rates to meet target performance goals.
+These cost parameters act as “tuning knobs,” which allow managers to adjust them to generate Pareto curves and balance congestion with overflow rates to meet target performance goals; see more in Appendix 12.3.
 We extend this model to ten-pool and twenty-pool systems.
@@
 Patients can overflow within their hospital or to the other hospital.
-Overflow costs are set at $B=(B_{1},B_{2},B_{3},B_{4},B_{5},B_{6},B_{7},B_{8})=(25,30,35,49,40,45,50,55)$, with lower costs for within-hospital overflows $(B_{1},\dots,B_{4})$ and higher costs for cross-hospital overflows $(B_{5},\dots,B_{8})$.
-The baseline unit holding costs remain the same, i.e., $C=6$ for regular and $C=7$ for VIP patients.
+Overflow costs are set at $B=(B_{1},\dots,B_{8})=(25,30,35,49,40,45,50,55)$, with lower costs for within-hospital overflows $(B_{1},\dots,B_{4})$ and higher costs for cross-hospital overflows $(B_{5},\dots,B_{8})$.
+The unit holding costs remain the same.
 […]
@@
 Total - 1260 280 0.889
-Table 2: Parameter settings for five-pool, ten-pool and twenty-pool models.
+Table 2: Parameter settings for five-pool, ten-pool and twenty-pool models; the latter two are formed by clustering and proportionally scaling the five-pool model in arrivals and capacities.
 6.2 Five-pool System: Comparing Performance with ADP
@@
 Furthermore, in the simpler two-pool systems where the exact optimal policy is known, PPO achieves near-optimal results (within 3% difference).
-More results are available in Appendix 12.
 Eight-epoch; Unbalanced Eight-epoch; Balanced
 B=(15,25) B=(30,35) B=(40,45) B=(15,25) B=(30,35) B=(40,45)
-ADP 152.68$\pm$0.76 196.31$\pm$0.92 250.39$\pm$0.45 132.46$\pm$0.32 193.03$\pm$0.62 248.48$\pm$0.71
-PPO 153.17$\pm$0.34 202.37$\pm$0.65 260.87$\pm$0.40 134.59$\pm$0.82 196.78$\pm$1.06 258.32$\pm$0.98
-Table 3: Five-pool system.
+[…]
+[…]
+Table 3: Five-pool system: long-run average costs.
 We set $N_{1}=60,N_{2}=64,N_{3}=67,N_{4}=N_{5}=62$ in unbalanced setting and $N_{1}=\cdots=N_{5}=63$ in balanced setting.
@@
 In larger systems with more than five pools, the ADP method becomes infeasible due to the need to exhaustively search over the combinatorially large action space.
-Therefore, we compare the performance of the PPO algorithm against three benchmark policies introduced in Dai and Shi (2019): (a) the complete-overflow policy, where overflow assignments are made as many as possible at each decision epoch; (b) the midnight policy, allowing overflow assignments only at midnight each day; and (c) the empirical policy, which reflects current practice by allowing overflow assignments only during nighttime (7 PM to 7 AM the next day).
-Figure 4 shows the performance comparison in ten-pool and twenty-pool systems, where PPO significantly outperforms all three naive policies.
+Therefore, we compare the performance of the PPO algorithm against three benchmark policies introduced in Dai and Shi (2019): (a) the complete-overflow policy, where overflow assignments are allowed at each decision epoch; (b) the midnight policy, allowing overflow assignments only at midnight each day; and (c) the empirical policy, which reflects current practice by allowing overflow assignments only during nighttime (7 pm to 7 am the next day).
+In all three policies, when overflow assignments are allowed, waiting patients are assigned one by one: first to their preferred overflow ward if feasible, and then to a randomly chosen secondary ward, until no waiting patients remain or no idle beds are available.
+Figure 4 shows the performance comparison in ten and twenty-pool systems, where PPO significantly outperforms all benchmark policies.
 In the ten-pool system, PPO achieves an average cost of 300, approximately 23% lower than the best-performing benchmark (the empirical policy with an average cost of 390).
@@
 The computational time is about 3.5 hours per iteration for the ten-pool system and about 8 hours per iteration for the twenty-pool systems.
-Note these are training times - once the policy networks are trained, the execution of the optimal policy is almost instantaneous in real-time deployment.
+Note these are training times; once the policy networks are trained, the execution of the policy is almost instantaneous in real-time deployment.
 These results highlight the scalability and effectiveness of our PPO algorithm in large-scale systems.
@@
 These results highlight the importance of adapting a general-purpose algorithm to the specific problem domain for optimal performance and computational efficiency.
+Table 4 summarizes our ablation study and the relative impact of each tailored component on performance and computation time.
+Configuration Performance Change Time Change
+Full algorithm
+Remove PPO (atomic + greedy search) $\downarrow\downarrow$ $\uparrow\uparrow$
+Remove tailored policy network (alternative NN designs) $\downarrow\downarrow$ $\sim$
+Remove tailored basis (basis in Dai and Shi (2019)) $\downarrow$ $\sim$
+Remove LSTD (NN critic) $\downarrow$ $\uparrow$
+Remove batching (sequential updates, various orders) $\sim$ $\uparrow\uparrow$
+Fewer epochs $\downarrow$ $\downarrow$
+Fewer simulation days $\sim$ $\downarrow$
+Table 4: Ablation study summary.
+For each variant, the term in parentheses indicates the replacement mechanism.
+$\uparrow\uparrow$” or $\downarrow\downarrow$” indicate large increase or decrease in performance and computation time, compared with the full algorithm (integrating all elements: Atomic + PPO + Tailored Policy Network + Tailored LSTD + Batching); $\uparrow$” or $\downarrow$” moderate increase or decrease; and $\sim$” negligible change.
+Is PPO necessary?
+Row 2 of Table 4 examines the effect of removing PPO and replacing it with a conventional value-based method (ADP) paired with atomic actions.
+In this setup, policy updates are performed via greedy search over all possible atomic actions.
+While atomic actions reduce the action space and make updates tractable, combining them with greedy search introduces significant instability due to the well-known “chattering” phenomenon (Bertsekas and others 2011).
+The average cost varying between 259.24 and 282.77, more than 27% increase from our PPO algorithm (see Appendix 12.2.2 for details).
+The chattering arises from the compounding effect of (i) the greedy policy updates (ii) the imperfect approximation of the value function, even with carefully chosen basis functions.
+These challenges are especially pronounced in long-run average cost settings, where small misestimations in value can lead to oscillatory behavior across iterations.
+Moreover, the atomic setup, which evaluates actions at an individual (rather than system) level, further exacerbates the chattering since each action’s marginal impact is highly sensitive to estimation noise.
+In contrast, pairing atomic actions with PPO mitigates these issues via incremental policy updates (as formalized in Proposition 1) and updating actions simultaneously via the system-level policy ratio (10).
+Together, these results highlight that the combination of atomic actions with PPO is essential in our infinite-horizon, long-run average context.
 Policy network design.
-We compare the performance of three policy NN designs in the ten-pool system across various neural network widths, depths, and training settings; see Figures 5-6.
+We compare the performance of three policy NN designs in the ten-pool system across various neural network widths, depths, and training settings; Row 3 of Table 4 summarizes the impact; Figures 5-6 show the detailed results.
 The figures plot the percentage improvement in average cost achieved by the PPO policy (after 10 iterations) compared to the best-performing benchmark, the empirical policy, thus, a higher percentage indicates greater improvement (more desirable).
@@
 (b) Impact of training epochsWhen training data is very small (1,000 simulation days per actor), to achieve good performance within fewer iterations, we use the no-overflow policy as the initial policy (instead of complete-sharing in the baseline) and set the training epochs to 30 (instead of 15 in the baseline).
-Figure 5: Trade-off between performance and training days and epochs.
-Performance is measured as the % improvement in average cost between the PPO policy and the best-performing benchmark (the empirical policy).
+Figure 5: Impact of training days and epochs. Performance is measured as the % improvement in average cost between the PPO policy and the best-performing benchmark (the empirical policy).
 (a) Impact of NN depth
 (b) Impact of NN width
-Figure 6: Trade-off between performance and NNs’ depth and width.
+Figure 6: Impact of NNs’ depth and width.
 Performance is measured as the % improvement in average cost between the PPO policy and the best-performing benchmark (the empirical policy).
@@
 Other tailored designs.
-Table 4 compares the average costs from using our proposed basis function against those used in Dai and Shi (2019), with varying amounts of simulation data.
-The table shows that our proposed basis functions are much less sensitive to the number of simulation days per actor.
-While the results with a larger dataset (10,000 days) are similar between the two sets of basis functions, the advantage of our proposed one becomes more evident with fewer simulation days (1,000 days) – the more practical choice for real-world implementation of the algorithm.
-Moreover, the results in Appendix 12.2 show that our proposed basis functions consistently obtains more accurate approximation among the top 25% most frequently visited states.
-Batching is another key design for improving time efficiency, as discussed in Section 5.3.
-We implement the original atomic decision process following Feng et al.
-(2021) for our problem and compare it with our batching setup.
-The batching setup delivers similar performance to the original but reduces computational time by approximately 50% (e.g., from two hours to one hour); see Appendix 12.3.
+Rows 4-5 of Table 4 report the impact of replacing our tailored LSTD with (a) using basis functions in Dai and Shi (2019) and (b) using a NN-based value network, respectively.
+Table 5 further compares the average costs of using our proposed basis functions against those in Dai and Shi (2019), under different amounts of simulation data.
+The results show that our tailored basis functions are much less sensitive to the number of simulation days: while performance is similar with a large dataset (10,000 days), the advantage of our proposed functions becomes more evident with fewer simulation days (1,000 days), which is more practical for real-world implementation.
+Using the NN-based value network, on the other hand, shows worse final performance and requires extra time for training.
+This is likely due to higher variance and less stable value estimates for training the NN.
+See Appendix 12.2.3 for detailed comparison.
+Row 6 of Table 4 compares the batching approximation and the original sequential atomic updates.
+The batching setup substantially reduces computational time ($\sim 50\%$) while maintaining similar performance.
 These findings show the value of our tailored designs in enhancing the time efficiency of the PPO algorithm.
-That is, by adapting to the queueing structure under the randomized atomic action setup, these tailored designs reduce sensitivity to the amount of simulation data.
 This not only speeds up the algorithm significantly but also makes it more practical for real-world applications where large-scale simulation data may be limited.
 Basis function 1k days per actor 10k days per actor
-Proposed 311.75$\pm$3.74 302.01$\pm$1.28
-Dai and Shi (2019) 325.71$\pm$4.07 307.71$\pm$1.31
+[…]
+[…]
 Table 5: Comparing the average costs from PPO algorithm using different basis functions.
-Impact of other parameters.
-We present in Appendix 12.1 a detailed performance comparison under various combinations of hyper-parameters.
+We present in Appendix 12.1 a detailed performance comparison under various combinations of hyper-parameters (remaining rows of Table 4).
 The key takeaway remains consistent with our earlier discussion: our tailored PPO design captures the most salient features of the problem and is of first-order importance in reducing long-run average costs and improving computational efficiency.
 While tuning other hyper-parameters, such as the number of simulation days, training epochs, and policy update iterations, can further enhance performance, their impact is relatively marginal.
-From the comparison results, we recommend increasing training epochs before expanding simulation data or iterations to maximize the use of existing data.
-Additionally, we identify two key techniques that further improve the algorithm’s efficiency:
-  1. 1.
-    Reuse of simulation data: Since PPO allows only small policy changes in each iteration, simulation data from a few previous iterations can be reused to reduce variance in estimating the value functions in the current iteration without introducing significant bias.
-This improves sample efficiency.
-Our experiments suggest that reusing data from the previous three iterations achieve the best results; see Table 12 in Appendix 12.3.
-  2. 2.
-    Adaptive clip size: A larger clip size can speed up policy improvement by allowing more substantial changes early in the training.
-However, it may hinder convergence to a stable policy later.
-Our experiments recommend starting with a larger clip size (e.g., 0.5) and reducing it as the policy update progresses; see Table 13 in Appendix 12.3.
-6.5 Practical Implementation
-A key strength of our approach is its domain adaptation: leveraging queueing structures and problem-specific characteristics to develop more effective and efficient algorithms.
-By integrating a tailored policy network, customized basis functions, and a batching setup, our PPO algorithm requires significantly less simulation data than typical reinforcement learning or ADP-based approaches.
-This not only accelerates training but also makes the algorithm more practical in environments with limited data or computational resources.
-The most important insights from our case study findings lies here as well, that is, the domain-aware adaptation is more critical to algorithm performance than fine-tuning neural network parameters in the original general-purposed version, which can be ad hoc and lack systematic guidance.
-For practical implementation, since our algorithm is model-based, the hospital first needs to calibrate the queueing system to reflect specific patient flows, arrival rates, and discharge patterns.
+6.5 Practical Implementation and Policy Interpretability
+In this section, we discuss how to implement our approach in practice, examine the policy interpretability since this is key for future implementation, and summarize actionable insights.
+Practical implementation.
+Since our algorithm is model-based, the hospital first needs to calibrate the queueing system to reflect specific patient flows, arrival rates, and discharge patterns.
 Such calibration has been successfully applied in various hospital settings, e.g, Bertsimas and Pauphilet (2024).
@@
 By adjusting these values, managers can explore trade-offs and generate Pareto-optimal solutions to meet their desired operational goals.
-For instance, increasing overflow costs could reduce the amount of overflow, or vice versa.
+For instance, increasing overflow costs could reduce the amount of overflow, or vice versa; see, e.g., Appendix 12.3.
 Note that the PPO algorithm employs a randomized policy that provides action probabilities for patient routing decisions.
-Decision-makers can use these probabilities, $\pi(f|s)$, to guide overflow strategies, such as selecting the most likely action or using de-randomization techniques (e.g., adding a final layer to the policy network for deterministic choices).
+Decision-makers can use these probabilities, $\pi(f\mid s)$, to guide overflow strategies, such as selecting the most likely action or using de-randomization techniques (e.g., adding a final layer to the policy network for deterministic choices).
 The randomized policy also offers flexibility for managers to adjust routing decisions in real time.
@@
 Ultimately, decisions regarding which individual patients to overflow will depend on the judgment of hospital managers, hospitalists, and charge nurses; our algorithm supports, rather than replaces, the overflow decision-making.
+Policy Interpretability.
+Using visualization and explainable AI techniques, we analyze the policy generated by our PPO algorithm and extract actionable insights for hospital managers.
+First, the learned policy aligns with managerial intuition in several key areas: it overflows more patients when source wards are congested and/or destination wards are underutilized; it is less aggressive with overflow during the day, anticipating discharges, and more proactive in the evening when discharges are rare; and it dynamically balances patient preferences with system-wide congestion levels when selecting overflow destinations.
+Beyond these intuitive behaviors, the analysis also uncovers more subtle, non-obvious network-level effects.
+For example, the policy proactively reroutes patients to mitigate anticipated future congestion in interconnected wards; jointly adjusts overflow decisions across patient classes sharing downstream resources, accounting for the cascading “chain effects” of the network structure.
+These results demonstrate that the PPO policy not only aligns with operational intuition but also leverages system-level complexity to achieve performance gains that would be difficult to replicate via manual decision rules.
+Full visualizations and detailed analyses are provided in Appendix 13.
+Practical and Algorithmic Insights.
+A key strength of our approach lies in its domain adaptation—leveraging problem-specific characteristics to develop more effective and efficient algorithms tailored in the hospital overflow context.
+By integrating a tailored policy network, customized basis functions, and a batching setup, our PPO algorithm requires significantly less simulation data than typical reinforcement learning or ADP-based approaches.
+This not only accelerates training but also makes the method practical in environments with limited data or computational resources.
+The case study highlights that domain-aware design choices are much more critical to algorithm performance than ad-hoc neural network tuning.
+Our findings offer a systematic path for adapting general-purpose RL frameworks to complex operational systems.
+Beyond algorithmic contributions, our policy visualization analysis also provides actionable heuristics for hospital managers, for example, overflow intensity should vary by time of day, and decisions should depend on the relative congestion of both source and target units rather than follow rigid rules.
+While such principles can support better decision-making in the absence of sophisticated analytics, our study also highlights the limitations.
+Human-devised heuristics often fall short in capturing the anticipatory dynamics and interconnected network effects of large hospital systems.
+Scalable, data-driven tools like ours can systematically learn and adapt to these complexities, going beyond what ad hoc rules can achieve.
 7 Conclusion
-This paper presents a scalable and efficient PPO algorithm for managing patient overflow in large hospital systems.
 Overflow management becomes increasingly complex as the number of hospital wards and patient classes increases, imposing challenges that traditional methods like ADP struggle to handle at scale.
-Our approach addresses this scalability issue by leveraging atomic actions and a partially-shared policy network.
-Additionally, our tailored designs for the hospital-specific queueing structures allow the training of our algorithm efficient and reduce the need for extensive simulation data, which overcomes a common limitation of reinforcement learning algorithms.
+The approach developed in this paper addresses the scalability issue by leveraging atomic actions and a partially-shared policy network.
+Additionally, our tailored designs for the hospital overflow context allow the training of our algorithm efficient and reduce the need for extensive simulation data, which overcomes a common limitation of reinforcement learning algorithms.
 The case study results demonstrate that our PPO algorithm matches or outperforms existing benchmarks in systems with up to twenty hospital wards.
+The atomic action + PPO framework is particularly effective in settings with the following structural features: (i) combinatorial action spaces involving multi-customer assignment decisions to multiple service pools; (ii) intertemporal congestion effects, where current decisions impact future system state and cost through queue dynamics; (iii) periodic, long-run optimization with arrival and departure patterns follow strong time-of-day or day-of-week cycles, which requires policies to adapt dynamically over time while still optimizing long-run performance.
+While we apply the method to inpatient overflow management, similar structures arise in other healthcare and service systems such as dynamic nurse assignment (Sun et al.
+2025), call centers or customer support systems, emergency responder dispatching, ride-hailing (Feng et al.
+2021), and long-term care admissions (Dong et al.
+2025).
+In these contexts, our approach offers a scalable alternative to full enumeration or problem-specific heuristics.
 There are several promising directions for future research building on the framework developed in this paper.
@@
 Incorporating a human-in-the-loop component will be crucial for balancing algorithmic recommendations with human decision-making to make the decision-support adaptable to the complexities of hospital operations.
-[…]